### v1

# Ab initio spin Hamiltonians and magnetism of Ce and Yb triangular-lattice compounds

Abstract
We calculate the crystal-field splitting, ground-state Kramers doublet and intersite exchange interactions within the ground-state doublet manifold using an ab initio Hubbard-I based approach for a representative set of Ce and Yb triangular-lattice compounds. These include the putative quantum spin liquids (QSL) RbCeO2 and YbZn2GaO5 and the antiferromagnets KCeO2 and KCeS2. The calculated nearest-neighbor (NN) couplings are antiferromagnetic and exhibit noticeable anisotropy. The next-nearest-neighbor (NNN) couplings are ferromagnetic in the Ce systems and dominated by classical dipole–dipole interactions in the Yb case. Solving the resulting effective spin-1/2 models by exact diagonalization up to $N=36$ sites, we predict ordered magnetic ground states for all systems, including the two QSL candidates. We explore the phase space of an anisotropic NN + isotropic NNN triangular-lattice model finding that a significant antiferromagnetic NNN coupling is required to stabilize QSL phases, while the NN exchange anisotropy is detrimental to them. Our findings highlight a possibly important role of deviations from the perfect triangular model – like atomic disorder – in real triangular-lattice materials.
Introduction. Triangular lattice spin-$1/2$ antiferromagnets (AFM), with their inherent geometric frustration and quantum fluctuations, stand as prime candidates for harboring exotic states of quantum matter, such as quantum spin liquids (QSL) and topologically ordered phases. While the nearest-neighbor Heisenberg model on the triangular lattice has been shown to stabilize a $120^{\circ}$ magnetically ordered Néel ground-state [1, 2, 3, 4], the inclusion of additional interactions has revealed a variety of both magnetically ordered and quantum disordered phases. Notably, enhanced second nearest-neighbor interactions can induce a quantum paramagnetic state governed by the physics of the gapless $U(1)$ Dirac spin liquid (DSL) [5, 6, 7], where distinct proposals have suggested a direct emergence of the DSL [8, 9, 10], a valence bond solid [11], a gapped $\mathbf{Z}_{2}$ spin liquids [12, 13, 14, 15] and chiral spin liquids [16, 17, 18], which are understood to be descendants of the parent DSL order [6, 7, 19, 20]. While the precise nature of the paramagnetic regime is still under debate, it has become evident that the physics of the DSL and its descendants serves as an organizing principle of the phase diagrams of triangular lattice antiferromagnets [11].
Geometrically perfect quasi-2D triangular lattices of magnetic ions are realized in numerous layered rare-earth oxides and chalcogenides. In particular, Kramers 3+ ions of RR=Ce, Yb, Nd etc. in systems like delafossites $ARX_{2}$ (where $A$ is an alkali metal, XX=O, S, Se) [21, 22, 23, 24, 25, 26, 27], heptatantalates RRTa7O19 [28], YbMgGaO4 [29, 30, 31, 32], and YbZn2GaO5 [33] have recently attracted a lot of interest. The crystal-field (CF) splitting in these systems is typically much larger than rather weak superexchange between the well localized 4ff shells of rare-earth ions. Consequently, admixture of the excited CF levels by superexchange can be neglected with the lowest Kramers doublet considered as an effective spin-1/2. However, strong spin-orbit (SO) entanglement in the 4ff ground state multiplet in conjunction with the weaker CF effect results in anisotropic charge and magnetic density of the Kramers states, which leads, in turn, to anisotropic superexchange and $g$-factors. Hence, the intersite coupling anisotropy that has been shown to strongly affect QSL formation is naturally present in those compounds. A gapless QSL ground state has been suggested on the basis of intensive experimental investigation for NaYbO2 [21], several Yb delafossite selenides and sulfides [34, 35] as well as for YbZn2GaO5 [33].
Low-energy Hamiltonians for those rare-earth triangular materials remain to date quite uncertain, in particular, due to lack of single crystals for delafossites [33]. It appears to be difficult to precisely extract small two-site anisotropies from magnon dispersions in an ordered state induced by applied field, hence, isotropic NN and NNN interactions are often assumed in the analysis [36] possible admixture of excited CF levels by magnetic field further complicating the analysis. Theoretically, ab initio calculations of exchange interactions in Yb delafossites on the basis of DFT electronic structure have so far predicted too small NNN couplings compared to experimental expectations [37, 36].
KCeO2 $0.85|5/2;\mp 1/2\rangle\pm 0.47|5/2;\pm 5/2\rangle\pm 0.18|7/2;\mp 1/2\rangle$
KCeS2 $0.90|5/2;\mp 1/2\rangle\pm 0.44|5/2;\pm 5/2\rangle-0.05|7/2;\mp 7/2\rangle$
RbCeO2 $0.87|5/2;\mp 1/2\rangle\pm 0.44|5/2;\pm 5/2\rangle\pm 0.17|7/2;\mp 1/2\rangle$
YZGO $0.73|7/2;\mp 7/2\rangle-0.50|7/2;\pm 5/2\rangle\mp 0.47|7/2;\mp 1/2\rangle$
Table 1: Ground-state Kramers doublet wavefunctions in the total-moment eigenstates $|JM\rangle$ basis. In the case of Ce oxides, the admixture of excited $J=7/2$ SO multiplet to the GS WV involve additional contributions apart from the largest one shown, as expressed by the dots at the end (see the SM [56] for the complete expressions).
In this work, we tackle the problem of evaluating realistic spin Hamiltonians for triangular rare-earth systems using DFT+dynamical mean-field theory [38, 39, 40] treating strongly localized rare-earth 4ff states in a quasiatomic (Hubbard-I) approximation [41] to obtain the high-temperature paramagnetic electronic structure, as well as, the 4ff-shell CF splitting and ground-state Kramers doublet. Intersite exchange interactions (IEI) are evaluated from the same paramagnetic electronic structure using the force-theorem in Hubbard-I (FT-HI) method [42]. In contrast to the previous ab initio calculations deriving IEI in triangular rare-earth systems from DFT electronic structure with metallic 4ff bands [37], within our framework the IEI are obtained from realistic self-consistent DFT+HI electronic structure of the target compounds with localized 4ff states. A significant impact of DFT+DMFT charge-density self-consistency on the electronic structure and CF splitting is well documented [43, 44, 45]. We apply this framework to derive realistic spin Hamiltonians for a representative set of Ce- and Yb- triangular compounds, namely: KCeO2, KCeS2, RbCeO2, and YbZn2GaO5 (YZGO). These materials are either experimentally inferred to exhibit conventional magnetic orders (KCeO2 [23] and KCeS2 [22, 24]), a putative DSL (YZGO [33, 46]) or remain to be investigated in detail (RbCeO2, for which no order was found in a preliminary study [25], but a full characterization has not been carried out). We then study the resulting Hamiltonians through exact diagonalization (ED) calculations [47, 48], obtaining numerically exact ground- and low-energy eigenstates resolved by the lattice space group on clusters with sizes up to $36$ spins. This, then, allows us to analyze in detail the nature of the magnetic ground-state (GS) spin-spin correlations for each of the considered compounds.
Ab initio method. We employ a charge self-consistent implementation [49, 50, 51] of DFT+DMFT based on the Wien2k linearized augmented-plane-wave (LAPW) full-potential code [52] and ”TRIQS” library [53]. All calculations were carried out using the reported experimental lattice structures [23, 24, 25, 33]. The IEI are calculated by FT-HI approach [42] using the MagInt code [54] analogously to previous applications of this approach to correlated magnetic insulators, see Ref. [55] for a review. The FT-HI method includes all kinetic exchange contributions due to virtual hopping of 4ff electrons. See Supplemental Material (SM) for more details of our ab initio calculations [56].
Figure 1: Calculated CF level splitting. The thick red lines are calculated CF levels; the dashed blue lines are CF excitations measured in INS from Refs. [23, 22, 25, 33]. Note that three excitations are seen in INS spectra for the Ce compounds.
KCeO2 KCeS2 RbCeO2 YZGO
$g_{ab}$ This work 1.43 2.00 1.54 3.24
Refs. 2.00a 1.67b, 2.47c 1.46d 3.44e
$g_{c}$ This work 0.17 -0.10 0.30 3.04
Refs. 0.29a 0.58b,0.65c 0.01d 3.04e
Table 2: Calculated values for the in-plane ($g_{ab}$) and out-of-plane ($g_{c}$) gyromagnetic tensor components. Other values are obtained by electron spin resonance, Ref. [23] (a); quantum chemistry calculations, Ref. [22] (b), Ref. [24] (c), Ref. [25] (d); magnetization measurements, Ref. [33] (e).
Crystal-field splitting and ground state. DFT+HI calculations converge to the expected insulating state with $f^{1}$ and $f^{13}$ GS occupancy of the Ce and Yb 4ff shells, respectively (see the SM for the calculated DFT+HI spectral functions). The average calculated SO coupling $\lambda$=87 meV for the Ce compounds leading to the SO gap between the GS $J=5/2$ and excited $J=7/2$ multiplets of 0.30 eV, with at most 1.5% spread among the 3 compounds. The value $\lambda=$0.39 eV with the SO splitting of 1.37 eV is obtained for Yb in YZGO. The calculated CF splitting are shown in Fig. 1 and the corresponding GS Kramers doublets are listed in Table 1 (see the SM for a complete list CF wavefunctions). We find a particularly large total CF splitting – about 60% of the SO gap – for the Ce oxides. Correspondingly, there is a non-negligible admixture $\approx 0.17|7/2;\pm 1/2\rangle$ of the SO excited multiplet to their GS Kramers doublets. The CF splitting in KCeS2 is almost 3 times smaller leading to the corresponding reduction of the $J=7/2$ admixture. No noticeable multiplet mixing is found for YZGO. One may notice the dominating contribution of $|M=\pm 1/2\rangle$ to the GS of the Ce systems corresponding to an in-plane magnetic anisotropy. In contrast, one may expect an out-of-plane anisotropy in YZGO due to the largest $|M=\pm 7/2\rangle$ contribution to its doublet. These expectations are confirmed by our results on the $g$-tensor (Table 2) extracted from the calculated magnetic moment operators $M_{\alpha}=g_{\alpha}S^{\alpha}$, $\alpha=x,y,z$ within the GS Kramers doublet. Here we introduced the pseudo-spin $S=1/2$ to label the states within the GS doublet and the corresponding spin operators $S^{\alpha}$. One sees that the in-plane components $g_{ab}=g_{x}(=g_{y})$ largely dominate in all Ce systems. In KCeS2, the out-of-plane component even becomes negative as a result of a near-perfect cancellation between the J=J=5/2 states, with the moment induced by a J=J=7/2 admixture. The magnitude of $g$ in YZGO is significantly larger, with a rather small out-of-plane anisotropy.
The experimental determination of the CF splitting and ground state in the Ce compounds under consideration is complicated by the appearance of an additional mode in their inelastic neutron scattering spectra [21, 25, 22]. A $J=5/2$ multiplet split into three Kramers doublet should result in two measurable CF excitations; instead, three excitations are observed. The lowest SO excited $J=7/2$ level is predicted by our calculations to be at about 300 meV agreeing with experiment [23]. Secondary oxide phases [24], vibronic splitting [57, 58] or more complicated multisite many-body effects [59] have been considered as possible origins of the additional mode, which is also observed in layered Ce intermetallics [59] and Ce pyrochlores [60], but it remains so far an unsolved puzzle. Taking this complication into account, our predictions for the CF levels energies in all compounds under consideration are in good agreement with experiment (Fig. 1), albeit with some systematic underestimation of the splitting by the present theory. Similarly, a good agreement is found for the $g$-tensor (Table 2) in comparison with previous experimental estimates and quantum chemistry calculations. We note that no signal was observed in the electronic spin resonance for the out-of-plane direction in KCeS2 suggesting $g_{c}\ll g_{ab}$ [22].
Effective magnetic Hamiltonian. Having obtained the GS Kramers doublet in all systems, we subsequently calculated the IEI within this manifold using the FT-HI method. Besides IEI, we also included the classical intersite dipole-dipole interaction, which can be important given rather small magnitudes of anisotropic IEI in the target materials, evaluated using the ab initio gyromagnetic tensors (Table 2).
The symmetry of the triangular rare-earth layer constrains the most general coupling within the $S=1/2$ manifold for the NN $ij$ bond along $\mathbf{r}_{0}^{\mathrm{NN}}=[100]||x$ to [29, 61, 62]:
$H_{ij}$ $=\mathbf{S}_{i}^{T}\hat{J}\mathbf{S}_{j}=J(\Delta S_{i}^{z}S_{j}^{z}+S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y})+$ (1)
$2J_{\pm\pm}(S_{i}^{x}S_{j}^{x}-S_{i}^{y}S_{j}^{y})+J_{z\pm}(S_{i}^{z}S_{j}^{y}+S_{i}^{y}S_{j}^{z}),$
where $\mathbf{S}^{T}_{i}=[S^{x}_{i},S^{y}_{i},S^{z}_{i}]$, $\Delta$ is the diagonal XXZ anisotropy, $J_{\pm\pm}$ is the diagonal XY anisotropy and $J_{z\pm}$ are off-diagonal terms. The coupling for the NNN bond along $\mathbf{r}_{0}^{\mathrm{NNN}}=[010]||y$ takes the same form of Eq. (1); we will use the prime to distinguish the NNN terms ($^{\prime},J±±J_{\pm\pm}^{\prime}$ etc.). The coupling for other NN and NNN bonds follows from the symmetry of the relevant point group that is generated by the threefold rotation about the $z$ axis, $C_{3}$, two-fold rotation $C_{2}$ around each axis $\mathbf{a}_{1},\mathbf{a}_{2},\mathbf{a}_{3}$, of the triangular lattice, and inversion $I$. The spin operator vector $\mathbf{S}$ transforms as a usual 3D vector upon proper rotations [63, 62]. By applying the $C_{3}$ and $C_{3}^{2}$ rotations about the $z$ axis, one obtains the couplings for the corresponding bonds as $\hat{J}_{\mathbf{r}}=\mathcal{R}\hat{J}_{\mathbf{r}_{0}}\mathcal{R}^{T}$, where $\mathbf{r}=\mathcal{R}\mathbf{r}_{0}$ and $\mathcal{R}$ is the corresponding rotation matrix. The couplings for the bonds related by inversion are identical, since inversion does not affect the spins. The symmetry of Eq. (1) is not imposed in FT-HI calculations, in which all matrix elements of (1) are evaluated independently, nevertheless it appears in the calculated IEI.
KCeO2 KCeS2 RbCeO2 YZGO
$J$ (meV) 0.51 0.021 0.36 0.25
$\Delta$ 1.03 1.90 1.05 1.03
$J_{\pm\pm}/J$ 0 0.18 0 -0.055
$J_{z\pm}/J$ 0.10 -1.45 -0.121 0.039
$J^{\prime}/J$ -0.055 -0.023 -0.063 0.003
$\Delta^{\prime}$ 0.70 -0.64 0.67 5.96
$J_{\pm\pm}^{\prime}/J^{\prime}$ 0.009 -0.62 -0.002 2.62
$J_{z\pm}^{\prime}/J^{\prime}$ 0.043 1.00 -0.040 0.073
$J^{{}^{\mathrm{int}}}/J$ 0.011 0.016 0.012
$\theta$ (K) -8.5 -0.43ab -6.2 -3.9ab,-1.9c
$\theta_{\mathrm{Exp}}$ (K) -7.7 -2.8ab -5.1 -5.2ab,-3.8c
Table 3: Calculated intersite coupling including the IEI and dipole-dipole contributions. Experimental Curie-Weiss temperatures $\theta_{\mathrm{Exp}}$ extracted from out-of-plane (c), in-plane (ab), or polycrystalline data are from Refs. [23, 22, 25, 33].
The resulting intersite couplings are listed in Table 3. All systems exhibit AFM NN $J$, as expected; we find $\Delta\gtrsim 1$ everywhere. The NNN interactions are ferroic in all Ce systems and antiferroic in YZGO. There is an essential difference between the Ce and Yb systems in the nature of NNN interactions: in the former, the dipole-dipole interactions are negligible compared to the exchange. In YZGO, due to small NNN IEI and large $g$ factors, the anisotropic dipole-dipole contribution dominates over an almost isotropic NNN exchange. As a result, the total NNN couplings in YZGO are very anisotropic ($\Delta\approx 6$) with $J_{xx}^{\prime}(=J^{\prime}+2J^{\prime}_{\pm\pm})=0.020J$, $J_{yy}^{\prime}(=J^{\prime}-2J^{\prime}_{\pm\pm})=-0.013J$, and $J_{yy}^{\prime}=\Delta J^{\prime}=0.019J$. Hence, though the NNN interactions in YZGO are small, they are not as negligible as may seem from the $J^{\prime}/J$ ratio listed in Table 3. The effect of a strong NNN anisotropy on the phase stability on the triangular lattice has not been studied so far. Finally, the coupling $J^{\mathrm{int}}$ for the shortest interlayer bond is about 1%-1.5% of $J$ in all oxide systems (we could not extract it in KCeS2 due to the general smallness of IEI in this system).
Figure 2: Spin structure factors and low-energy spectrum for the YZGO (a,b) and KCeS2 (c,d) compounds in a $36$-site cluster. The isotropic structure factor $\mathcal{S}\left(\bm{k}\right)$ in (a) is peaked at the $K$ points, indicating $120^{\circ}$ order, consistent with low-energy excitations at $K_{0}.\text{A}$ and $K_{1}.\text{B}$. The peak at $M_{2}$ of $\mathcal{S}^{yy}(\bm{k})$ in (c) indicates stripe-$\perp$ order, consistent with low-energy excitations at the $M$ points in (d).
To assess the overall agreement in magnitude with experiments, we calculated high-temperature susceptibility by solving (1) in mean-field in an applied field using the McPhase package [64] and an in-house module. To compare with polycrystalline data we averaged the calculated susceptibility as $(2\chi_{ab}+\chi_{c})/3$. For all oxide systems we find quite good agreement of the calculated Curie-Weiss temperatures $\theta$ with available experimental data (Table 3). In the case of KCeS2, theoretical intersite couplings are severely underestimated. This is also apparent from comparison of our result with the estimates of Ref. [65] that were obtained by quantum chemistry calculations and subsequently adjusted to agree with experimental INS. 4ff kinetic exchange captured by FT-HI does not, apparently, give the leading contribution to the IEI in this sulfide.
Low-energy spectra and ground-state correlations. We now investigate the GS phase diagram of the effective magnetic Hamiltonians introduced in Eq. (1) with ED using the XDiag library [48]. We obtain symmetry-resolved excitation spectra employing translational symmetries and the $I$ lattice inversion symmetry without spin rotation on cluster sizes $N=12,32,36$, see the SM [56] for the employed lattice geometries. The irreducible representations (irreps) are represented by $\bm{k}.\rho$, where $\bm{k}$ labels the momentum and $\rho$ labels the inversion irrep, i.e. either $\rho=\text{A}$ for the even irrep or $\rho=\text{B}$ for odd irrep. We note that the additional point group symmetries are not employed, as an implementation of combined lattice-spin rotations is not readily accessible. We also obtained the GS static spin structure factors,
$\mathcal{S}^{\alpha\alpha}\left(\bm{k}\right)=\dfrac{1}{N}\sum_{n,m}e^{i\bm{k}\cdot\left(\bm{r}_{n}-\bm{r}_{m}\right)}\left\langle S^{\alpha}_{n}\cdot S^{\alpha}_{m}\right\rangle,$ (2)
where $\alpha=x,y,z$, $N$ denotes the number of lattice sites and $\bm{r}_{n}$ the position of the $n^{\rm th}$ spin. Furthermore, we denote by,
$\mathcal{S}\left(\bm{k}\right)=\mathcal{S}^{xx}\left(\bm{k}\right)+\mathcal{S}^{yy}\left(\bm{k}\right)+\mathcal{S}^{zz}\left(\bm{k}\right),$ (3)
the isotropic spin structure factor.
In panels $a)$ and $b)$ of Fig. 2, we show $\mathcal{S}\left(\bm{k}\right)$ and the low-lying energy spectrum for the model parameters of YZGO (see table 3) on the $N=36$ cluster. Both the organization of the spectrum and the spin structure factor are consistent with the system realizing the $120^{\circ}$ AFM phase. The ground state transforms under the $\Gamma.\text{A}$ irrep, while the first excited states—nearly degenerate—belong to the $\Gamma.\text{B}$ and $K.\text{A}$ irreps. These symmetry quantum numbers match the expected signatures of the $120^{\circ}$ AFM phase [66, 67], see the SM [56] for further details on how these quantum number are predicted. A key signature of a DSL is a gapless singlet mode at the $X_{0}$ point in the first Brillouin zone (FBZ), which constitutes the singlet monopole excitations [19]. However, the excitation energy at $X_{0}$ is significantly larger compared to the putative DSL of the $J_{1}$-$J_{2}$ model [11], rendering this scenario unlikely. Furthermore, the static spin structure factor exhibits a pronounced peak at crystal momentum $K$ within the FBZ, in agreement with the known ordering wave vector of the $120^{\circ}$ AFM state [68, 69]. Analogously, we observe spin correlations and a low-energy spectrum consistent with a $120^{\circ}$ state for the effective Hamiltonians of the KCeO2 and RbCeO2 compounds, see the SM [56].
Figure 3: $a)$ and $d)$ - Approximate phase diagram of the effective magnetic Hamiltonian, inferred from the organization of quantum numbers in the low-energy spectrum. The dark red region indicates the $120^{\circ}$ antiferromagnetic phase, dark blue the stripy-$\parallel$ order, pale purple the stripy-$\perp$ order, and orange the regime where the first non-zero momentum excitation is in the $X_{1}.\text{A}$ irrep, which suggests a possible DSL phase. A typical classical ordering for each stripe phase is shown in the sketch inside panel d). The remaining panels show the static spin structure factor at selected momenta in the FBZ (see inset of panel $a$). Panels $b$ and $e$ display $\mathcal{S}(\mathbf{k})$ at $Z$, while $c$ and $f$ show the average $\tilde{\mathcal{S}}(M)$ over the three inequivalent $M$ points. In panel $b)$ we locate the parameters compatible with the YZGO system. Parameters: $\Delta=1.03J$, $J^{\prime}=0.02J$ in a)a)$c)$, and $J_{z\pm}=0$ in d)d)$f)$.
The spectrum and correlations for the KCeS2 parameters, on the other hand, indicate a collinear stripy antiferromagnetic GS, as shown in panels $c)$ and $d)$ of Fig. 2. The low energy spectra reveals an approximate GS degeneracy between the $\Gamma$.A, and $M.\text{A}$ irreps, consistent with the stripy Néel state [67]. Accordingly, we observe a pronounced peak in $\mathcal{S}^{yy}(\bm{k})$ and $\mathcal{S}^{zz}(\bm{k})$ at momentum $M_{2}$We observe the peak only at one of the nonequivalent $M$ wave vectors, since the $C_{3}$ rotational symmetry of the Hamiltonian is slightly broken by numerical rounding in the coupling constants obtained from the ab-initio calculations.. As noticed previously [71], the symmetries of Eq. 1 allow for two distinct stripe orders, differing by how they transform under the combined lattice-spin rotation around the triangular lattice axis directions. The first, we here call stripe-$\parallel$ (stripe-parallel), aligns the spins along the direction of the stripe. It is invariant under the $C_{2}$ axis rotations as the spins align with the axis of rotation. The stripe-$\perp$ (stripe-perpendicular) state aligns the spins perpendicular to the stripe direction, and thus transforms non-trivially under the $C_{2}$ axis rotation. A representative configuration of each stripe type is shown in Fig. 3d). In general, this state is not invariant under the $C_{2}$ axis rotation. Thus, the (in)variance under this symmetry constitutes the distinguishing feature between the two states. We prefer to refer to these states as stripe-$\parallel$ and stripe-$\perp$ as we think this more precisely captures the essence of these phases than their previously used names stripe-$x$ and stripe-$yz$, respectively.
As the ground state stripe stabilized is oriented along the $x$-direction of the lattice, the peak in the $\mathcal{S}^{yy}(\bm{k})$ and $\mathcal{S}^{zz}(\bm{k})$ at $M_{2}$ indicates the realization of the stripe-$\perp$ state, see the SM for further details. Remarkably, this is in full agreement with experiments on KCeS2 [24].
[…]
We obtain an approximate phase diagram, shown in panel $a)$ and $d)$ of Fig. 3, based on the quantum numbers of the spectrum, which we show in the supplementary material. We begin by fixing $J^{\prime}=0.02J$ (a value numerically compatible with the largest matrix elements of the NNN coupling $\hat{J^{\prime}}$ in YZGO; see Table 3) and study the phase diagram as a function of the diagonal XY anisotropy $J_{\pm\pm}$ and the off-diagonal XY coupling $J_{z\pm}$. This corresponds to the panels (a,b,c) in the upper row of Fig. 3. In this case, the phase diagram consists of three distinct magnetically ordered states, namely, a $120^{\circ}$ AFM region, a stripe-$\parallel$ phase, and a stripe-$\perp$ phase, which correspond to the dark red, dark blue, and pale purple regions shown in panel $a)$, respectively.
As on the 32-site cluster, the $K$ momentum is not resolved, the structure factor peaks at nearby wave vectors $Z_{0}$ and $Z_{1}$ in the $120^{\circ}$ AFM phase, as shown in panel (b) of Fig. 3. This reflects the frustrated realization of $120^{\circ}$ AFM order in this cluster. In contrast, the 12-site cluster resolves the $K$ points, showing a structure factor peak there consistent with the $120^{\circ}$ AFM order (see supplementary material). We observe in panel b) that the couplings determined for the YZGO system are within the $120^{\circ}$ AFM as expected.
Upon tuning both $J_{\pm\pm}$ and $J_{z\pm}$, $\mathcal{S}(\bm{k})$ develops a peak at the $M$ points, signaling the onset of the stripe phase, as shown in panels $c)$ and $d)$. From the phase diagram in panel (a), we find that along the $J_{z\pm}=0$ line, negative values of $J_{\pm\pm}/J$ favor a stripe-$\parallel$ GS, whereas positive values of $J_{\pm\pm}/J$ stabilize a stripe-$\perp$ state. The coupling $J_{z\pm}$ further extends the stripe-$\perp$ region into negative $J_{\pm\pm}/J$ values, which agrees with previous studies. In certain parameter regimes, the peak in $\mathcal{S}(\bm{k})$ appears in both the $M_{0}$ and $M_{1}$ wave vectors, while in others it shifts to the $M_{2}$ wave vector. This shift does not signify a transition between the two stripe phases, but instead arises from the geometry of the $32$-site cluster (being absent in the $12$-site cluster). As the $32$-site cluster does not have the $C_{3}$ rotation symmetry, the three $M$ points are inequivalent. As a result, the degeneracy among the six stripe configurations is lifted by the lattice geometry. To account for this, in panel (c) of Fig. 3, we display
$\tilde{\mathcal{S}}(M)=[\mathcal{S}(M_{0})+\mathcal{S}(M_{1})+\mathcal{S}(M_{2})]/3$ (4)
averaged over the three $M$ points. Distinguishing between the two stripe types requires examining additional spin-spin correlators, which is left to the SM.
We now set $J_{z\pm}=0$ and study the phase diagram as a function of the NNN coupling $J^{\prime}$ and $J_{\pm\pm}$. Based on the analysis of the quantum numbers of the lowest excited states, we map out the phase diagram identifying four distinct regions, as shown in panel $d)$ of Fig. 3. For $J^{\prime}/J\sim[0.0,0.05]J$, the GS is in an $120^{\circ}$ AFM state, which is further corroborated by the peak in $\mathcal{S}(\bm{k})$ at the $Z$ points (shown in panel $e)$). The $120^{\circ}$ AFM is stabilized up to values of $J_{\pm\pm}\sim 0.1J$. Increasing $J_{\pm\pm}$ promotes the development stripe order, whose precise nature depends on the sign of $J_{\pm\pm}/J$, as shown in panel $d)$. Negative values favor a stripe-$\parallel$ GS, whereas positive values stabilize a stripe-$\perp$. The stripe order is further corroborated by the strong intensity of $\mathcal{S}(\bm{k})$ at the $M$ points, as shown in the panel $f)$. The black and white lines indicating the phase boundaries in Fig. 3 are obtained from the position of level crossings in the excitation spectrum, see SM [56] for details.
In the intermediate regime, $J^{\prime}\sim[0.05,0.16]J$ and $J_{\pm\pm}\sim\pm 0.1J$, the first excited state outside the $\Gamma.\text{A}$ sector transforms according to the $X_{1}.\text{A}$ irrep, as shown in the supplementary material. This parameter range corresponds to the orange region in the phase diagram (panel $d)$). In the context of triangular-lattice antiferromagnets, the appearance of an excitation at momentum $X$ is particularly noteworthy, as it has been interpreted as a signature of a DSL [19, 20]. Specifically, Ref. [19] predicted that the singlet-monopole excitation of the DSL carries momentum $X$ and appears in the low-energy spectrum of the $J_{1}$$J_{2}$ triangular-lattice antiferromagnet in the spin-liquid regime [11]. An alternative explanation for such a low-lying excitation is a 12-site valence-bond solid, for which consistent correlations have also been observed in the spin-liquid regime [11]. In this parameter window, the static spin structure factor exhibits nearly the same intensity at the $Z$ and $M$ points. We further find that this region is stabilized only for small values of $J_{\pm\pm}\sim\pm 0.1J$, while larger values drive the system towards the stripe phase.
Conclusions – We have derived from first principles the crystal-field splitting, Kramer’s GS and intersite exchange interactions within the GS manifold for a set of rare-earth triangular materials, including two systems (KCeO2 and KCeS2) that order magnetically, as well as, two systems (RbCeO2 and YZGO) that have been reported as putative QSLs. The calculated crystal-field splitting and gyromagnetic $g$-factors are in good agreement with experiment. The mean-field Curie-Weiss temperature evaluated from the resulting low-energy Hamiltonian including both the calculated intersite exchange and classical dipole-dipole magnetic interaction agrees well with experiment for all three oxide systems but is severely underestimated for KCeS2.
Solving the resulting low-energy Hamiltonian with ED, we find evidence for the $120^{\circ}$ AFM order in all systems apart from KCeS2, where a stripe-$\perp$ AFM state is instead stabilized, consistent with experimental observations. We find strong evidence for a magnetic $120^{\circ}$ order in the two DSL candidates. From the extended phase diagram of their Hamiltonian, we concluded that a stronger antiferromagnetic NNN coupling is required to realize the DSL phases in comparison to those present in the compounds.
We notice that a very recent neutron-scattering study [76] reports significant site disorder in YZGO as evidenced by broadening of crystal-field excitations and small random crystalline distortions. The effect of site disorder was discussed in the context of a well-studied similar system, YbMgGaO4, where the resulting randomization of intersite exchange was predicted to mimic a QSL state [32]. Our results together with the observations of Ref. [76] suggest that the same phenomenon may be at play in YZGO.
Our calculated intersite exchange interactions include all $f$-$f$ kinetic exchange contributions (in particular, superexchange). However, in rare-earth semiconductors there can be another type of $f$-$f$ intersite exchange generated through 4ff-5dd hopping and the inter-shell Hund’s rule interaction [77]. This could become important in the case of very weak superexchange like in KCeS2, where our calculations apparently reproduce the structure of intersite exchange but not its magnitude. Further studies to evaluate the importance of such exchange mechanisms in triangular rare-earth compounds are worthwhile.
Acknowledgments – The exact diagonalization calculations have been performed using the XDiag library [48]. L.V.P. is grateful to Sidhartha Dash, Dario Fiore Mosca, and Antoine Georges for useful discussions and to the CPHT computer team for support. A.W. acknowledges support by the German Research Foundation (DFG) through the Emmy Noether program (Grant No. 509755282).
Supplemental Material to ‘Ab initio spin Hamiltonians and magnetism of Ce and Yb triangular-lattice compounds’
Appendix A Ab initio methodology and calculational details
[…]
In the case of Ce delafossites, the Ce 4ff KS band manifold is separated from the ligand $p$-band by a wide gap and it only slightly overlaps with the Ce 5dd band. The lower boundary of the energy window is thus chosen to be within the $pf$ gap; the upper boundary is placed at 1.5 eV above the KS Fermi level $E_{F}^{KS}$, right above the top of the 4ff band.
[…]
In order to correct for this problem in YZGO, we thus apply, within Wien2k, a uniform shift to the 4ff band position to keep its centerweight position aligned to that in pure DFT (0.9 eV below $E_{F}^{KS}$) in the course of DFT+HI iterations. To disentangle 4ff from other contributions to hybridized 2pp-4ff bands, we extend the Wannier basis to include ligand O 2pp and itinerant Zn 3dd states. It has been shown [79] that employing a narrow energy window to construct 4ff Wannier orbitals is crucial to capture the hybridization contribution to CF. Since this narrow window excludes most of O 2pp and Zn 3dd states, we employ two energy windows. The 4ff Yb orbitals are formed from applying the initial projection to the bands in the range [-1.7:1.7] eV around the centerweight of KS 4ff band, which fully encloses this band. The O 2pp and Zn 3dd Wannier orbitals are formed from projection in the large window of [-9.5:1.5] eV around $E_{F}^{KS}$ that fully includes both corresponding band manifolds. The resulting full set of Wannier orbitals including Yb 4ff, O 2pp and Zn 3dd is then orthonormalized using the standard prescription of the projective construction approach [49] to obtain a proper Wannier basis.
We parametrize the 4ff-shell rotationally-invariant on-site Coulomb repulsion by $U=F^{0}$ and the Hund’s rule coupling $J_{H}$ using the values U=U=6(8) eV and $J_{H}=$0.7(1.1) eV for the Ce(Yb) 4ff shells, respectively. The values of $U$ and $J_{H}$ together with the standard assumptions on the ratios of the Slater parameters ($F^{4}/F^{2}$=0.668, $F^{6}/F^{2}$=0.45 [80]) fully determine the vertex. The chosen $J_{H}$ values are consistent with experimental data on RE 3+ ions in insulators [81], the values of $U$ are within the generally accepted range and increasing from Ce to Yb in agreement with theoretical estimates for the RE series [82, 83]. We employed the fully localized limit for the DMFT double-counting correction, which was calculated using the nominal $f^{1}$ ($f^{13}$) occupancy in the Ce (Yb) cases as shown to be appropriate for the DFT+HI method [43]. To avoid the DFT self-interaction error impacting the CF parameters we employ averaging over the ground-state multiplet in our self-consistent calculations in accordance with Ref. [45]. Namely, the Boltzmann weights for all levels within the ground-state multiplet $^{2}F_{5/2}$ ($^{2}F_{7/2}$) of the Ce (Yb) 4ff shell are averaged in the course of self-consistent DFT+HI calculations.
Having converged DFT+HI calculations, we compute the intersite exchange interactions between the CF ground-state Kramer’s doublets using the FT-HI method. This method extract the low-energy exchange Hamiltonian from the paramagnetic electronic structure by evaluating the response of DFT+DMFT grand potential to small fluctuations around the paramagnetic state on two neighboring sites, see Ref. [42] for details. To form a proper (pseudo-)spin-1/2, the phases of the Kramer’s doublet states are chosen to have the proper symmetry under the time reversal $T$, $T|+1/2\rangle=|-1/2\rangle$.
Appendix B DFT+Hubbard-I paramagnetic electronic structure
In Fig. S1 we display the electronic structure of the paramagnetic phase in Ce delafossites (exemplified by KCeO2) and YZGO calculated by the DFT+Hubbard-I approach. In the case of KCeO2, the Ce 4ff lower Hubbard band is located in the wide semiconducting gap between O 2pp and Ce 5dd. In the case of the Yb compounds, the 4ff lower Hubbard band is within the O 2pp band and also overalaps with the filled Zn 3dd states, while the 4ff upper Hubbard band is located about 3.5 eV above the bottom of the conduction band (the latter is mainly formed by Yb 5dd states hybridized with O 2pp).
Figure S1: DFT+Hubbard-I spectral functions of KCeO2 (top panel) and YbZn2GaO5 (bottom panel).
Appendix C Crystal field levels and wavefunctions
[…]
Energy (meV) Wavefunction
KCeO2
0 $0.853|5/2;\mp 1/2\rangle\pm 0.470|5/2;\pm 5/2\rangle\pm 0.177|7/2;\mp 1/2\rangle-0.134|7/2;\mp 7/2\rangle+0.041|7/2;\pm 5/2\rangle$
117 $0.993|5/2;\pm 3/2\rangle\mp 0.116|7/2;\pm 3/2\rangle+0.039|7/2;\mp 3/2\rangle$
167 $0.828|5/2;\pm 5/2\rangle\mp 0.484|5/2;\mp 1/2\rangle+0.243|7/2;\mp 1/2\rangle\pm 0.145|7/2;\mp 7/2\rangle$
KCeS2
0 $0.895|5/2;\mp 1/2\rangle\pm 0.444|5/2;\pm 5/2\rangle-0.045|7/2;\mp 7/2\rangle$
38 $0.998|5/2;\pm 3/2\rangle\mp 0.055|7/2;\pm 3/2\rangle+0.032|7/2;\mp 3/2\rangle$
68 $0.890|5/2;\pm 5/2\rangle\mp 0.441|5/2;\mp 1/2\rangle+0.087|7/2;\mp 1/2\rangle\pm 0.074|7/2;\mp 7/2\rangle$
RbCeO2
0 $0.872|5/2;\mp 1/2\rangle\pm 0.439|5/2;\pm 5/2\rangle\pm 0.169|7/2;\mp 1/2\rangle-0.132|7/2;\mp 7/2\rangle+0.039|7/2;\pm 5/2\rangle$
109 $0.993|5/2;\pm 3/2\rangle\mp 0.111|7/2;\pm 3/2\rangle+0.040|7/2;\mp 3/2\rangle$
164 $0.851|5/2;\pm 5/2\rangle\mp 0.453|5/2;\mp 1/2\rangle+0.229|7/2;\mp 1/2\rangle\pm 0.131|7/2;\mp 7/2\rangle$
YbZn2GaO5
0 $0.725|7/2;\mp 7/2\rangle-0.501|7/2;\pm 5/2\rangle\mp 0.471|7/2;\mp 1/2\rangle$
21 $0.831|7/2;\mp 5/2\rangle+0.555|7/2;\pm 7/2\rangle\pm 0.033|7/2;\pm 1/2\rangle$
54 $|7/2;\pm 3/2\rangle$
84 $0.881|7/2;\mp 1/2\rangle\pm 0.408|7/2;\mp 7/2\rangle\mp 0.238|7/2;\pm 5/2\rangle$
Table 1: Calculated crystal-field energies and wavefunctions in all four compounds.
Figure S2: Geometry of the simulation clusters with (a) $N=12$, (b) $N=32$, and (c) $N=36$ spins. The shaded area indicates the simulation box, with periodic boundary directions taken along the edges.
Appendix D Simulation clusters used in the ED calculations
The simulation clusters that we use in our calculations are the $N=12,32$ and $36$ sites shown in Fig. S2. The shaded region corresponds to the simulation box with periodic boundary conditions taken along each of the edges. The $N=12$ and $36$ site clusters are compatible with the full point-group symmetries of the infinite triangular lattice, the full dihedral group of order 12 ($D_{6}$), whereas the $N=32$ site cluster is not due to its specific geometry and its symmetries correspond to the full dihedral group of order 4 ($D_{2}$). As discussed in the main text, the $12$ and $36$ site clusters are compatible with all the symmetries of the anisotropic Hamiltonian, while the $32$ site is not.
Appendix E Quantum Numbers of the symmetry broken states
In the thermodynamic limit, the ground state of magnetically ordered systems spontaneously breaks the symmetries of the Hamiltonian. This breaking is reflected in the organization of quantum numbers in the low-energy spectrum, forming the so-called Anderson Tower of States. Here we predict their quantum numbers using group representation theory, following Ref. [67].
We perform this analysis for two types of stripe-ordered states, stripe-$\parallel$ and stripe-$\perp$. The corresponding classical product states are shown in panels $a)$ and $b)$ of Fig. S3, respectively. The two stripe orders differ in how they transform under combined lattice-spin $C_{2}$ rotation around the lattice axis. As a result, the quantum numbers in panels $a)$ and $b)$ differ in their point-group irrep: the stripe-$\parallel$ states transform under the trivial irrep, while the stripe-$\perp$ states transform under the odd irrep. The multiplicities of the irreps in the degenerate ground-state manifold are calculated considering the point-group $D_{3d}$. However, in our ED calculations we only use translational and $C_{2}$ lattice rotation symmetries (ignoring spin rotations). Consequently, we do not resolve the quantum numbers of the entire symmetry group shown in panel $a)$ and $b)$. Under the used symmetries in the ED calculations, the two types of stripe order have identical signatures in the spectra, as shown in panel $c)$. Nevertheless, we can identify the transition between the two stripe types through a level crossing in the spectrum, specifically of the states that transform under the irreps with momentum $M$.
We have also carried out this analysis for the $120^{\circ}$ AFM state, again using only translational and $C_{2}$ lattice rotation symmetries. The corresponding quantum numbers are shown in panel $d)$.
Figure S3: Scheme of the six degenerate classical ground states for $a)$ stripes-$\parallel$ and $b)$ stripes-$\perp$, together with the multiplicities ($n$) of the corresponding irreducible representations under the full $D_{3d}$ point-group symmetry of the Hamiltonian. c)d)c)-d) Multiplicities of the irreducible representations for both stripe-$\parallel$ and stripe-$\perp$ (table $c)$) and for the $120^{\circ}$ AFM phase (table $d)$) when the Hamiltonian possesses only the $C_{2}$ lattice symmetry.
Appendix F ED results effective Hamiltonians of KCeO2 and RbCeO2
Figure S4: $a)$ Low-energy spectrum, $b)$ $\mathcal{S}(\bm{k})$ and $c)$ spin correlations $\langle\bm{S}_{\bm{r}_{j}}\cdot\bm{S}_{\bm{r}_{0}}\rangle$ in the GS of the $36$-site cluster effective magnetic Hamiltonians for the different compounds considered in this work (each row corresponds to a different compound). The spin correlations are computed with respect to the reference site $r_{0}$ marked by the green cross.
Here, we discuss the low-energy spectra and GS correlations on the effective Hamiltonians derived for the RbCeO2 and KCeO2 materials (consult Eq. (1) and Tab. 3). These results are shown in the third and fourth row of Fig. S4. For comparison, we show in the same figure the results for the YZGO and KCeS2 compounds that were previously discussed in the main text.
Similarly to the YZGO material, we observe an organization of the low-energy excitations and spin correlations pointing to the $120^{\circ}$ AFM order. First, in the low-energy excitations, we observe that the GS lies in the $\Gamma.A$ irrep, with the first excited state corresponding to the $\Gamma.\text{B}$ being almost degenerate with a pair of excitations at momentum $K$, as shown in panel $a)$ for both RbCeO2 and KCeO2 materials. These excitations in this symmetry sectors are compatible with those expected for the $120^{\circ}$ AFM phase (see Fig. S3 d)). Secondly, $\mathcal{S}(\bm{k})$ peaks at the $K$ wave-vectors (shown in panels $b)$) . The real-space spin correlations [panels $c$] exhibit the sign-changing pattern characteristic of the three-sublattice structure of the $120^{\circ}$ order, reflecting the alternating spin orientations imposed by the triangular lattice geometry and frustrated antiferromagnetic interactions. This pattern is found in all compounds except KCeS2, which instead shows GS correlations characteristic of stripe order. Along the $\hat{x}$ direction the spins align ferromagnetically, while antiferromagnetic correlations remain along $\hat{y}$, consistent with the strong peak of $\mathcal{S}(\bm{k})$ at momentum $M_{2}$. To determine whether the system realizes the stripe-$\parallel$ or stripe-$\perp$ phase, we analyze $\mathcal{S}^{xx}(\bm{k})$, $\mathcal{S}^{yy}(\bm{k})$, and $\mathcal{S}^{zz}(\bm{k})$, shown in panels a)a)$c)$ of Fig. S5. All three components peak at $M_{2}$, confirming stripe order oriented along $100]\parallel x$. However, the much stronger correlations in $\mathcal{S}^{yy}(M_{2})$ and $\mathcal{S}^{zz}(M_{2})$ indicate that the GS realizes the stripe-$\perp$ configuration.
Figure S5: a) $\mathcal{S}^{xx}(\bm{k})$, b) $\mathcal{S}^{yy}(\bm{k})$ and c) $\mathcal{S}^{zz}(\bm{k})$ in the 36-site cluster for the effective Hamiltonian GS of the KCeS2 compound.
Appendix G Low-energy spectrum for the $N=32$ cluster and Further Results on the Spin Structure Factor
Here, we present a detailed analysis of the low-energy spectrum of the effective magnetic Hamiltonian on the $N=32$ site cluster used in the main text to construct the extended phase diagram. In addition, we discuss the static spin structure factor, which allow us to distinguish between the stripe-$\parallel$ and stripe-$\perp$ phases.
We begin with the spectra. Fig. S6 shows the dependence of the low-energy levels on different symmetry sectors as a function of the model parameters. Panel (a) displays the dependence on $J_{\pm\pm}$ at fixed $J_{z\pm}=0.173J$, taken from Fig. 3(a), while panel (b) shows the dependence on $J^{\prime}$ at fixed $J_{\pm\pm}=-0.0488J$, taken from Fig. 3(d). The different magnetic phases can be identified from the organization of the quantum numbers. The $120^{\circ}$ AFM region is signaled by a first excited state in the $\Gamma.\text{A}$ irrep together with two degenerate excited states in the $Z_{0}.\text{A}$ and $Z_{1}.\text{A}$ irreps. The transition to either stripe phase is marked by the closure of the gap in the $\Gamma.\text{A}$ irrep, corresponding to the phase boundaries in the main text. This criterion agrees remarkably well with the magnetic order inferred from the spin structure factors, as shown in the main text. Deep in the stripe phase, one also expects three degenerate states in the $M_{0}.\text{A}$, $M_{1}.\text{A}$, and $M_{2}.\text{A}$ irreps, becoming quasi-degenerate with the ground state in the $\Gamma.\text{A}$ irrep, as clearly seen in Fig. S6(a). The putative DSL regime is indicated by a first excited state with non-zero momentum $X$. This regime is shown in Fig. S6(b), where the $X_{1}.\text{A}$ level becomes the lowest excitation with a finite momentum as $J^{\prime}/J$ increases. We emphasize that this criterion alone does not establish the existence of the DSL phase nor determine its precise boundaries. We represent the transition to and from this region by a dashed line to emphasize this point.
As discussed above, the two types of stripe order cannot be distinguished solely from the quantum numbers of the spectra within the symmetries accessible to our ED calculations. However, they can be differentiated by analyzing the spin structure factor with spin components projected along the stripe direction and rotated by $90^{\circ}$ from it. Due to the symmetries of the $32$-site cluster, the six stripe configurations sketched in Fig. S3 are no longer degenerate. We find that in some regions of the phase diagram the ground state exhibits a characteristic wave vector at the $M_{2}$ point, while in others the peak appears at both $M_{0}$ and $M_{1}$ wave vectors. This distinction does not yet separate stripe-$\parallel$ from stripe-$\perp$; rather, it indicates which specific stripe arrangement (within each stripe type) is energetically preferred for this cluster geometry.
So for the parameters where the order is characterized by the $M_{2}$ wave vector, we inspect $\mathcal{S}^{xx}(\text{M}_{2})$ and $\mathcal{S}^{yy}(\text{M}_{2})$, shown in Figs. S6(c) and (d). In the stripe-$\parallel$ phase, the spins preferably align along the stripe direction (the $x$-axis), leading to $\mathcal{S}^{xx}(M_{2})>\mathcal{S}^{yy}(M_{2})$. In contrast, in the stripe-$\perp$ phase the spins are rotated by $90^{\circ}$, resulting in $\mathcal{S}^{yy}(M_{2})>\mathcal{S}^{xx}(M_{2})$. From this criterion, we identify the stripe-$\parallel$ region for $J_{\pm\pm}>0$ and the stripe-$\perp$ region for $J_{\pm\pm}<0$.
Nevertheless, there are regions within the stripe-ordered phases where neither $\mathcal{S}^{xx}(M_{2})$ nor $\mathcal{S}^{yy}(M_{2})$ show a peak. These correspond to regions where the GS develops stripe order with characteristic wave vectors at $M_{1}$ (or $M_{0}$). In this case, we analyze instead
$\mathcal{S}^{\parallel\bm{a}_{2}}\left(\text{M}_{1}\right)=\dfrac{1}{N}\sum_{i,j}\left\langle\left(\bm{S}_{i}\cdot\hat{\bm{a}}_{2}\right)\left(\bm{S}_{j}\cdot\hat{\bm{a}}_{2}\right)\right\rangle,\quad\mathcal{S}^{\perp\bm{a}_{2}}\left(\text{M}_{1}\right)=\dfrac{1}{N}\sum_{i,j}\left\langle\left(\bm{S}_{i}\cdot[1-\hat{\bm{a}}_{2}]\right)\left(\bm{S}_{j}\cdot[1-\hat{\bm{a}}_{2}]\right)\right\rangle,$ (S1)
where $\hat{\bm{a}}_{2}$ is the unit vector in the $\bm{a}_{2}$ lattice direction. These are shown in Figs. S6(e) and (f) as a function $J_{\pm\pm}/J$ and $J^{\prime}/J$. From these we obtain the same result: stripe-$\parallel$ is stabilized for $J_{\pm\pm}>0$, while stripe-$\perp$ is stabilized for $J_{\pm\pm}<0$. Moreover, the transition between them, occurring around $J_{\pm\pm}=0$, agrees remarkably well with the one identified from the crossing of the ground state and first excited state in the $M$-sector irreps. The behavior is identical when analyzing the dependence on $J_{\pm\pm}$ and $J_{z\pm}$ with $J^{\prime}$ fixed.
Figure S6: Low-energy spectra and static spin structure factors of the effective magnetic Hamiltonian for the 32-site cluster. $a)$ - Spectrum as a function of $J_{\pm\pm}/J$ for $J^{\prime}=0.022J$ and $J_{z\pm}=0.173J$. $b)$- Spectrum as a function of $J^{\prime}/J$ for $J_{\pm\pm}=-0.0488J$ and $J_{z\pm}=0$. The colored regions indicate phases: blue – stripe-$\parallel$, dark blue – stripe-$\perp$, red – $120^{\circ}$ AFM, pale red – putative DSL. c)f)c)–f) - Static spin structure factors: (c) along $x$ and (d) along $y$ at momenta $M_{2}$; (e) parallel and (f) perpendicular to $\bm{a}_{2}$ at momenta $M_{1}$.
Appendix H Low-energy spectrum and ground-state correlations for the $N=12$ cluster
In this section, we discuss the phase diagram obtained for the 12-site cluster, whose results support those presented in the main text for the 32-site cluster. In contrast to the 32-site cluster, the 12-site cluster resolves the natural ordering vectors for both the stripe order and the $120^{\circ}$ AFM order, specifically the $M$ and $K$ wave-vectors, as depicted in Fig. S7 $d)$. Moreover, the $X$ points are also resolved, which is important for the DSL phase.
Figure S7: a)-Approximate phase diagram as a function of $J_{\pm\pm}/J$ and $J_{z\pm}/J$ for different values of $J^{\prime}/J$ [i)–iii)]. The blue cross marks the approximate coupling values of some of the compounds considered in this work (we are using the absolute value of $J_{z\pm}$ see the main text). $\mathcal{S}(\bm{k})$ is shown at the ordering vectors $K_{0}$ [iv–vi] and $M$ [vii–ix], as well as for the $x$ [x–xii] and $y$ [xiii–xv] spin components at the $M_{2}$ wave-vector. (b) Low-energy spectra resolved by symmetry sectors as a function of $J_{\pm\pm}/J$ for $J_{z\pm}=0.20J$ [i)] and $J_{z\pm}=0.69J$ [ii)], with $J^{\prime}=0.02J$. (c) Sketch of the FBZ for the 12-site cluster.
Figure S8: $a)$ Approximate phase diagram based on the structure of the quantum numbers of the low-energy spectra as a function of $J_{\pm\pm}/J$ and $J^{\prime}/J$. The transition line between the two distinct stripe orders is obtained from the level crossing in the $M.\mathrm{A}$ irreps. $b)$ and $c)$ Low-energy spectra for $J_{\pm\pm}=0.0202J$ as a function of $J^{\prime}/J$ $b)$ and for $J^{\prime}=0.1054J$ as a function of $J_{\pm\pm}/J$ $c)$. The color scheme labeling the different irreps is the same as in Fig. S7. d)d)$f)$ Static spin structure factor for different as a function of $J_{\pm\pm}/J$ and $J^{\prime}/J$ evaluated at momentum $K$ (panel $d)$) and $M_{2}$ (panels $e)$ and $f)$). Other parameters: $N=12$ and $J_{z\pm}=0$.
We begin by studying the phase diagram as a function of $J_{\pm\pm}/J$ and $J_{z\pm}/J$ for fixed values of $J^{\prime}$ consistent with those obtained for the compounds considered in this work (see Tab. 3 in the main text). The phase diagram is shown in panel a), sub-panels i), ii), and iii), corresponding to $J^{\prime}=0.020J$, $J^{\prime}=-0.041J$, and $J^{\prime}=-0.062J$, respectively. The approximate phase boundaries are estimated from the quantum numbers of the low-energy spectrum. In the $120^{\circ}$ phase, the GS is unique transforming under the $\Gamma.\text{A}$ irrep, with low-energy excitations in the $\Gamma.\text{B}$, $K_{0}.\text{A}$ and $K_{1}.\text{A}$ irreps. We also observe an exactly degenerate first excited state also transforming under $\Gamma.\text{A}$ irrep. These exact degeneracy is due to the richer symmetry group than the one considered in our ED code. Upon tuning either $J_{\pm\pm}/J$ or $J_{z\pm}/J$, we observe a closure of the energy gap between the GS energy and the first excited state within the $\Gamma.{\rm A}$ irrep. The exact point where this gap vanishes is identified as the transition into the stripe phase. This transition is further confirmed by the decrease in intensity in $\mathcal{S}(K)$ after the gap closes, as shown in Fig. S7 a), panels iv)–vi). In the phase diagrams shown in panel a), we have located some of the compounds considered in this work. Their positions are determined by the calculated values of $^{\prime}/J, J±±/J$$_{\pm\pm}/J$, and the absolute value of $J_{z\pm}$ (magnetic phases with $J_{z\pm}<0$ can be mapped to those with $J_{z\pm}>0$ by applying a $\pi$-rotation to all spins in the system). We find their locations to be consistent with previous calculations, namely within the $120^{\circ}$ AFM state.
In the stripe-ordered phase, the ground state becomes doubly degenerate in $\Gamma.{\rm A}$, while the first excited states are degenerate and transform under the $M_{0}.{\rm A}$, $M_{1}.{\rm A}$, and $M_{2}.{\rm A}$ irreps. This also corroborates the onset of stripy magnetic order, as these quantum numbers match those predicted from the representation group theory calculations (see panel c) in Fig. S3). This is again corroborated by the static spin structure factors evaluated at the $M$ point, as shown in Fig.S7 a), panels vii)–ix).
To distinguish between the stripe-$\parallel$ and stripe-$\perp$,we examine $\mathcal{S}^{xx}(\bm{k})$ and $\mathcal{S}^{yy}(\bm{k})$. Both exhibit peaks at the wave vector $M_{2}$, but in different regions of the phase diagram: the $xx$ peak identifies the stripe-$\parallel$ phase, while the $yy$ peak signals the stripe-$\perp$ phase. The phase diagram obtained from them is consistent with the transition line inferred from the energy-level crossing at the $M$ points. This crossing reflects the transition between the two stripe orders, which carry distinct quantum numbers in the low-energy sector (although both remain located at momentum $M$). This crossing is clearly illustrated in Fig. S7c), panels i) and ii), where we show the energy spectra as functions of $J_{\pm\pm}/J$ for two representative values of $J_{z\pm}$: $J_{z\pm}\simeq 0.20J$ (i)) and $J_{z\pm}\simeq 0.69J$ (ii)).
We find that the $120^{\circ}$ order is suppressed as either $J_{\pm\pm}$ or $J_{z\pm}$ is increased, with negative values of $J^{\prime}/J$ enlarging this region in the phase diagram. By tuning $J_{\pm\pm}$ and/or $J_{z\pm}$, the GS instead exhibits stripe order. Similar to the 32-site cluster, we observe that for $J_{z\pm}=0$, negative $J_{\pm\pm}$ values favors the stripe-$\parallel$ phase, while positive values stabilize the stripe-$\perp$ phase. The stripe-$\perp$ phase is also supported for negative values of $J_{\pm\pm}/J$, when $J_{z\pm}\neq 0$. For all these parameters, we do not find the first excited state (with a non-zero momentum) in the $X.\mathrm{A}$ irrep, which therefore does not indicate the presence of a DSL phase.
In Fig. S8, we set $J_{z\pm}=0$ and study the phase diagram as a function of $J_{\pm\pm}/J$ and $J^{\prime}.J$. In panel (a), the phase boundaries are determined in the same way as for Fig. S7. For $J^{\prime}/J\lesssim 1\%$, the behavior matches that discussed previously: a $120^{\circ}$ AFM phase is stabilized for finite $J_{\pm\pm}/J$, which, upon increasing $|J_{\pm\pm}/J|$, gives way to a stripy phase. The nature of this phase depends on the sign of $J_{\pm\pm}/J$: for negative values it is stripe-$\parallel$, while for positive values it is stripe-$\perp$. This is corroborated by the static spin structure factors shown in panels d) to f): in the $120^{\circ}$ phase we find strong peaks at both $K$ points, whereas in the stripe-$\parallel$ phase the $S^{xx}$ component exhibits a pronounced signal along the $M_{2}$ vector, and in the stripe-$\perp$ phase the same wave-vector is dominant in the $S^{yy}$ component. The transition between the two stripe phases is again identified by locating the level crossing between the ground state and the first excited state within the $M.\text{A}$ irreps.
Increasing the strength of NNN coupling, $J^{\prime}/J$, leads to the appearance of first excitated state with non-zero momentum at the the $X_{1}.A$ and $X_{2}.A$ irrep. This is clearly seen in panels (b) and (c), where we plot the low-energy spectrum as a function of $J^{\prime}$ for $J_{\pm\pm}=0.02J$ and as a function of $J_{\pm\pm}$ along the line $J^{\prime}=0.1054J$, respectively. As discussed in the main text, this behavior pins the possible DSL region. From panel (a), we observe that the range of $J_{\pm\pm}/J$ supporting this region shrinks as $J^{\prime}$ increases. For sufficiently large values of either coupling, the system transitions into a stripy ordered phase whose orientation again depends on the sign of $J_{\pm\pm}/J$ (negative for stripe-$\parallel$, positive for stripe-$\perp$). Exactly at $J_{\pm\pm}=0$, the model recovers the SU(2) and U(1) spin-symmetries. These findings are consistent with the results for the 32-site cluster presented in the main text: while a possible DSL regime is identified, all ab initio estimates of $J^{\prime}/J$ lie below the required range.

### v2

@@
 A gapless QSL ground state has been suggested on the basis of intensive experimental investigation for NaYbO2 [21], several Yb delafossite selenides and sulfides [34, 35] as well as for YbZn2GaO5 [33].
-Low-energy Hamiltonians for those rare-earth triangular materials remain to date quite uncertain, in particular, due to lack of single crystals for delafossites [33].
-It appears to be difficult to precisely extract small two-site anisotropies from magnon dispersions in an ordered state induced by applied field, hence, isotropic NN and NNN interactions are often assumed in the analysis [36] possible admixture of excited CF levels by magnetic field further complicating the analysis.
+Low-energy Hamiltonians for those rare-earth triangular materials remain to date quite uncertain.
+In particular, due to lack of single crystals for majority of delafossites [33], two-site anisotropies were assessed experimentally only in few cases [36, 27].
+Hence, isotropic NN and NNN interactions are often assumed in the analysis [37], which is further complicated by possible admixture of excited CF levels by magnetic field.
 Theoretically, ab initio calculations of exchange interactions in Yb delafossites on the basis of DFT electronic structure have so far predicted too small NNN couplings compared to experimental expectations [38, 37].
-KCeO2 $0.85|5/2;\mp 1/2\rangle\pm 0.47|5/2;\pm 5/2\rangle\pm 0.18|7/2;\mp 1/2\rangle$
-KCeS2 $0.90|5/2;\mp 1/2\rangle\pm 0.44|5/2;\pm 5/2\rangle-0.05|7/2;\mp 7/2\rangle$
-RbCeO2 $0.87|5/2;\mp 1/2\rangle\pm 0.44|5/2;\pm 5/2\rangle\pm 0.17|7/2;\mp 1/2\rangle$
-YZGO $0.73|7/2;\mp 7/2\rangle-0.50|7/2;\pm 5/2\rangle\mp 0.47|7/2;\mp 1/2\rangle$
-Table 1: Ground-state Kramers doublet wavefunctions in the total-moment eigenstates $|JM\rangle$ basis.
-In the case of Ce oxides, the admixture of excited $J=7/2$ SO multiplet to the GS WV involve additional contributions apart from the largest one shown, as expressed by the dots at the end (see the SM [56] for the complete expressions).
-In this work, we tackle the problem of evaluating realistic spin Hamiltonians for triangular rare-earth systems using DFT+dynamical mean-field theory [38, 39, 40] treating strongly localized rare-earth 4ff states in a quasiatomic (Hubbard-I) approximation [41] to obtain the high-temperature paramagnetic electronic structure, as well as, the 4ff-shell CF splitting and ground-state Kramers doublet.
+In this work, we tackle the problem of evaluating realistic spin Hamiltonians for triangular rare-earth systems using DFT+dynamical mean-field theory (DMFT) [39, 40, 41] treating strongly localized rare-earth 4ff states in a quasiatomic (Hubbard-I) approximation [42] to obtain the high-temperature paramagnetic electronic structure, as well as, the 4ff-shell CF splitting and ground-state Kramers doublet.
 Intersite exchange interactions (IEI) are evaluated from the same paramagnetic electronic structure using the force-theorem in Hubbard-I (FT-HI) method [43].
@@
 This, then, allows us to analyze in detail the nature of the magnetic ground-state (GS) spin-spin correlations for each of the considered compounds.
+Previous works have already investigated certain regions of the parameter space for these type of Hamiltonians, through classical Monte Carlo studies [50, 51], ED [52] and notably using DMRG [32, 53, 54, 55].
 Ab initio method.
@@
 The FT-HI method includes all kinetic exchange contributions due to virtual hopping of 4ff electrons.
-See Supplemental Material (SM) for more details of our ab initio calculations [56].
+See the Supplemental Material (SM) [63] for more details on our ab initio calculations and the effect of charge-density self-consistency in DFT+HI.
 Figure 1: Calculated CF level splitting.
-The thick red lines are calculated CF levels; the dashed blue lines are CF excitations measured in INS from Refs.
+The thick red lines are calculated CF levels; the dashed blue lines are CF excitations measured in inelastic neutron scattering (INS) from Refs.
 [23, 22, 25, 33].
@@
 $g_{ab}$ This work 1.43 2.00 1.54 3.24
-Refs. 2.00a 1.67b, 2.47c 1.46d 3.44e
+Refs. 2.00a 1.67b, 2.47c 1.46d 3.04e
 $g_{c}$ This work 0.17 -0.10 0.30 3.04
-Refs. 0.29a 0.58b,0.65c 0.01d 3.04e
+Refs. 0.29a 0.58b,0.65c 0.01d 3.44e
 Table 1: Calculated values for the in-plane ($g_{ab}$) and out-of-plane ($g_{c}$) gyromagnetic tensor components.
-Other values are obtained by electron spin resonance, Ref.
-[23] (a); quantum chemistry calculations, Ref.
-[22] (b), Ref.
-[24] (c), Ref.
+Other values are obtained by electron spin resonance (ESR), Refs.
+[23] (a) ESR and quantum chemistry calculations (QCC) [22] (b); QCC, Ref.
+[24] (c); INS, Ref.
 [25] (d); magnetization measurements, Ref.
@@
 DFT+HI calculations converge to the expected insulating state with $f^{1}$ and $f^{13}$ GS occupancy of the Ce and Yb 4ff shells, respectively (see the SM for the calculated DFT+HI spectral functions).
-The average calculated SO coupling $\lambda$=87 meV for the Ce compounds leading to the SO gap between the GS $J=5/2$ and excited $J=7/2$ multiplets of 0.30 eV, with at most 1.5% spread among the 3 compounds.
-The value $\lambda=$0.39 eV with the SO splitting of 1.37 eV is obtained for Yb in YZGO.
 The calculated CF splitting are shown in Fig.
-1 and the corresponding GS Kramers doublets are listed in Table 1 (see the SM for a complete list CF wavefunctions).
-We find a particularly large total CF splitting – about 60% of the SO gap – for the Ce oxides.
+1 and the corresponding Kramers doublets are listed in the SM [63].
+We find a particularly large total CF splitting – about 60% of the SO gap (the latter is 0.3 eV in all Ce systems) – for the Ce oxides.
 Correspondingly, there is a non-negligible admixture $\approx 0.17|7/2;\pm 1/2\rangle$ of the SO excited multiplet to their GS Kramers doublets.
 The CF splitting in KCeS2 is almost 3 times smaller leading to the corresponding reduction of the $J=7/2$ admixture.
-No noticeable multiplet mixing is found for YZGO.
-One may notice the dominating contribution of $|M=\pm 1/2\rangle$ to the GS of the Ce systems corresponding to an in-plane magnetic anisotropy.
-In contrast, one may expect an out-of-plane anisotropy in YZGO due to the largest $|M=\pm 7/2\rangle$ contribution to its doublet.
-These expectations are confirmed by our results on the $g$-tensor (Table 2) extracted from the calculated magnetic moment operators $M_{\alpha}=g_{\alpha}S^{\alpha}$, $\alpha=x,y,z$ within the GS Kramers doublet.
+No noticeable multiplet mixing is found for YZGO due to its much larger SO gap of 1.37 eV.
+The dominating contribution of $|M=\pm 1/2\rangle$ to the GS of all Ce systems leads to an in-plane magnetic anisotropy.
+In contrast, the $|M=\pm 7/2\rangle$ states provide the largest contribution to the GS doublet of YZGO.
+The corresponding $g$-tensor extracted from the calculated magnetic moment operators $M_{\alpha}=g_{\alpha}S^{\alpha}$, $\alpha=x,y,z$ within the GS Kramers doublet is listed in Table 1.
 Here we introduced the pseudo-spin $S=1/2$ to label the states within the GS doublet and the corresponding spin operators $S^{\alpha}$.
@@
 In KCeS2, the out-of-plane component even becomes negative as a result of a near-perfect cancellation between the J=J=5/2 states, with the moment induced by a J=J=7/2 admixture.
-The magnitude of $g$ in YZGO is significantly larger, with a rather small out-of-plane anisotropy.
-The experimental determination of the CF splitting and ground state in the Ce compounds under consideration is complicated by the appearance of an additional mode in their inelastic neutron scattering spectra [21, 25, 22].
-A $J=5/2$ multiplet split into three Kramers doublet should result in two measurable CF excitations; instead, three excitations are observed.
-The lowest SO excited $J=7/2$ level is predicted by our calculations to be at about 300 meV agreeing with experiment [23].
-Secondary oxide phases [24], vibronic splitting [57, 58] or more complicated multisite many-body effects [59] have been considered as possible origins of the additional mode, which is also observed in layered Ce intermetallics [59] and Ce pyrochlores [60], but it remains so far an unsolved puzzle.
+The $g$-tensor in YZGO is significantly larger and quasi-isotropic.
+The experimental determination of the CF splitting and GS in the Ce compounds under consideration is complicated by the appearance of an additional, third, mode in their INS spectra [21, 25, 22], though only two CF excitations are expected in the case of a $J=5/2$ GS.
+This phenomenon, observed also in other Ce systems, remains so far without commonly accepted explanation, see Refs.
+[24, 64, 65].
 Taking this complication into account, our predictions for the CF levels energies in all compounds under consideration are in good agreement with experiment (Fig.
@@
 Effective magnetic Hamiltonian.
-Having obtained the GS Kramers doublet in all systems, we subsequently calculated the IEI within this manifold using the FT-HI method.
+The IEI within the GS Kramers manifold were calculated by the FT-HI method.
 Besides IEI, we also included the classical intersite dipole-dipole interaction, which can be important given rather small magnitudes of anisotropic IEI in the target materials, evaluated using the ab initio gyromagnetic tensors (Table 1).
@@
 (1); we will use the prime to distinguish the NNN terms ($^{\prime},J±±J_{\pm\pm}^{\prime}$ etc.).
-The coupling for other NN and NNN bonds follows from the symmetry of the relevant point group that is generated by the threefold rotation about the $z$ axis, $C_{3}$, two-fold rotation $C_{2}$ around each axis $\mathbf{a}_{1},\mathbf{a}_{2},\mathbf{a}_{3}$, of the triangular lattice, and inversion $I$.
-The spin operator vector $\mathbf{S}$ transforms as a usual 3D vector upon proper rotations [63, 62].
-By applying the $C_{3}$ and $C_{3}^{2}$ rotations about the $z$ axis, one obtains the couplings for the corresponding bonds as $\hat{J}_{\mathbf{r}}=\mathcal{R}\hat{J}_{\mathbf{r}_{0}}\mathcal{R}^{T}$, where $\mathbf{r}=\mathcal{R}\mathbf{r}_{0}$ and $\mathcal{R}$ is the corresponding rotation matrix.
-The couplings for the bonds related by inversion are identical, since inversion does not affect the spins.
-The symmetry of Eq.
-(1) is not imposed in FT-HI calculations, in which all matrix elements of (1) are evaluated independently, nevertheless it appears in the calculated IEI.
+The coupling for other NN and NNN bonds follows from the symmetry [67, 66], see SM [63] for details.
 KCeO2 KCeS2 RbCeO2 YZGO
@@
 Figure 2: Spin structure factors and low-energy spectrum for the YZGO ($a,b$) and KCeS2 ($c,d$) compounds in a $36$-site cluster.
-The isotropic structure factor $\mathcal{S}\left(\bm{k}\right)$ in (a) is peaked at the $K$ points, indicating $120^{\circ}$ order, consistent with low-energy excitations at $K_{0}.\text{A}$ and $K_{1}.\text{B}$.
-The peak at $M_{2}$ of $\mathcal{S}^{yy}(\bm{k})$ in (c) indicates stripe-$\perp$ order, consistent with low-energy excitations at the $M$ points in (d).
+The isotropic structure factor $\mathcal{S}\left(\bm{k}\right)$ in $a)$ is peaked at the $K$ points, indicating $120^{\circ}$ order, consistent with low-energy excitations at $K_{0}.\text{A}$ and $K_{1}.\text{B}$.
+The peak at $M_{2}$ of $\mathcal{S}^{yy}(\bm{k})$ in $c)$ indicates stripe-$\perp$ order, consistent with low-energy excitations at the $M$ points in $d)$.
 To assess the overall agreement in magnitude with experiments, we calculated high-temperature susceptibility by solving (1) in mean-field in an applied field using the McPhase package [68] and an in-house module.
-To compare with polycrystalline data we averaged the calculated susceptibility as $(2\chi_{ab}+\chi_{c})/3$.
-For all oxide systems we find quite good agreement of the calculated Curie-Weiss temperatures $\theta$ with available experimental data (Table 3).
+For all oxide systems we find quite good agreement of the calculated Curie-Weiss temperatures $\theta$, extracted from the averaged susceptibility $(2\chi_{ab}+\chi_{c})/3$, with available experimental data (Table 2).
 In the case of KCeS2, theoretical intersite couplings are severely underestimated.
@@
 We now investigate the GS phase diagram of the effective magnetic Hamiltonians introduced in Eq.
-(1) with ED using the XDiag library [48].
-We obtain symmetry-resolved excitation spectra employing translational symmetries and the $I$ lattice inversion symmetry without spin rotation on cluster sizes $N=12,32,36$, see the SM [56] for the employed lattice geometries.
+(1) with ED using the XDiag library [49, 70].
+We obtain symmetry-resolved excitation spectra employing translational symmetries and the lattice inversion symmetry without spin rotation on cluster sizes $N=12,32,36$, see the SM [63] for the employed lattice geometries.
 The irreducible representations (irreps) are represented by $\bm{k}.\rho$, where $\bm{k}$ labels the momentum and $\rho$ labels the inversion irrep, i.e.
@@
 $\mathcal{S}^{\alpha\alpha}\left(\bm{k}\right)=\dfrac{1}{N}\sum_{n,m}e^{i\bm{k}\cdot\left(\bm{r}_{n}-\bm{r}_{m}\right)}\left\langle S^{\alpha}_{n}\cdot S^{\alpha}_{m}\right\rangle,$ (2)
-where $\alpha=x,y,z$, $N$ denotes the number of lattice sites and $\bm{r}_{n}$ the position of the $n^{\rm th}$ spin. Furthermore, we denote by,
-$\mathcal{S}\left(\bm{k}\right)=\mathcal{S}^{xx}\left(\bm{k}\right)+\mathcal{S}^{yy}\left(\bm{k}\right)+\mathcal{S}^{zz}\left(\bm{k}\right),$ (3)
-the isotropic spin structure factor.
+where $\alpha=x,y,z$, $N$ denotes the number of lattice sites and $\bm{r}_{n}$ the position of the $n^{\rm th}$ spin.
+Furthermore, we denote by $\mathcal{S}\left(\bm{k}\right)=\sum_{\alpha}\mathcal{S}^{\alpha\alpha}\left(\bm{k}\right)$, the isotropic spin structure factor.
 In panels $a)$ and $b)$ of Fig.
@@
 Both the organization of the spectrum and the spin structure factor are consistent with the system realizing the $120^{\circ}$ AFM phase.
-The ground state transforms under the $\Gamma.\text{A}$ irrep, while the first excited states—nearly degenerate—belong to the $\Gamma.\text{B}$ and $K.\text{A}$ irreps.
-These symmetry quantum numbers match the expected signatures of the $120^{\circ}$ AFM phase [66, 67], see the SM [56] for further details on how these quantum number are predicted.
-A key signature of a DSL is a gapless singlet mode at the $X_{0}$ point in the first Brillouin zone (FBZ), which constitutes the singlet monopole excitations [19].
-However, the excitation energy at $X_{0}$ is significantly larger compared to the putative DSL of the $J_{1}$-$J_{2}$ model [11], rendering this scenario unlikely.
-Furthermore, the static spin structure factor exhibits a pronounced peak at crystal momentum $K$ within the FBZ, in agreement with the known ordering wave vector of the $120^{\circ}$ AFM state [68, 69].
-Analogously, we observe spin correlations and a low-energy spectrum consistent with a $120^{\circ}$ state for the effective Hamiltonians of the KCeO2 and RbCeO2 compounds, see the SM [56].
-Figure 3: $a)$ and $d)$ - Approximate phase diagram of the effective magnetic Hamiltonian, inferred from the organization of quantum numbers in the low-energy spectrum.
-The dark red region indicates the $120^{\circ}$ antiferromagnetic phase, dark blue the stripy-$\parallel$ order, pale purple the stripy-$\perp$ order, and orange the regime where the first non-zero momentum excitation is in the $X_{1}.\text{A}$ irrep, which suggests a possible DSL phase.
-A typical classical ordering for each stripe phase is shown in the sketch inside panel d).
-The remaining panels show the static spin structure factor at selected momenta in the FBZ (see inset of panel $a$).
-Panels $b$ and $e$ display $\mathcal{S}(\mathbf{k})$ at $Z$, while $c$ and $f$ show the average $\tilde{\mathcal{S}}(M)$ over the three inequivalent $M$ points.
-In panel $b)$ we locate the parameters compatible with the YZGO system.
+The GS transforms under the $\Gamma.\text{A}$ irrep, while the first excited states (nearly degenerate) belong to the $\Gamma.\text{B}$ and $K.\text{A}$ irreps.
+These symmetry quantum numbers match the expected signatures of the $120^{\circ}$ AFM phase [71, 72], see the SM [63] for further details on how these quantum numbers are predicted.
+A key signature of a DSL is a gapless singlet mode at the $X_{0}$ point in the first Brillouin zone (FBZ), corresponding to the singlet monopole excitations [19].
+Although in the present case we cannot identify these excitations as singlets due to the absence of spin SU(2) symmetry, they still occur at the same momentum.
+In contrast, the excitation energy at $X_{0}$ is significantly larger compared to the putative DSL of the $J_{1}$-$J_{2}$ model [11], rendering this scenario unlikely.
+Furthermore, $\mathcal{S}(\bm{k})$ exhibits a pronounced peak at crystal momentum $K$, in agreement with the known ordering wave vector of the $120^{\circ}$ AFM state [2, 73].
+Analogously, we also observe spin correlations and a low-energy spectrum consistent with a $120^{\circ}$ state for the effective Hamiltonians of the KCeO2 and RbCeO2 compounds, see the SM [63].
+Figure 3: $a)$ and $d)$ - Approximate phase diagram of the effective magnetic Hamiltonian, inferred from the quantum number organization of the low-energy spectrum.
+Dark red marks the $120^{\circ}$ AFM phase, dark blue the stripy-$\parallel$ order, pale purple the stripy-$\perp$ order, and orange the regime with first non-zero momentum excitation in the $X_{1}.\text{A}$ irrep, suggesting a possible DSL phase.
+A typical classical ordering for each stripe phase is shown in the sketch in $d)$.
+The remaining panels show the static spin structure factor at selected momenta in the FBZ (see inset in $a$)): panel $b)$ and $e)$ at $Z$, while $c)$ and $f)$ show the average over the three inequivalent $M$ points.
+In $b)$ we identify the parameters compatible with the YZGO system.
 Parameters: $\Delta=1.03J$, $J^{\prime}=0.02J$ in a)a)$c)$, and $J_{z\pm}=0$ in d)d)$f)$.
-The spectrum and correlations for the KCeS2 parameters, on the other hand, indicate a collinear stripy antiferromagnetic GS, as shown in panels $c)$ and $d)$ of Fig.
+On the other hand, the spectrum and correlations for the KCeS2 system indicate a collinear stripy antiferromagnetic GS, as shown in panels $c)$ and $d)$ of Fig.
 2.
-The low energy spectra reveals an approximate GS degeneracy between the $\Gamma$.A, and $M.\text{A}$ irreps, consistent with the stripy Néel state [67].
-Accordingly, we observe a pronounced peak in $\mathcal{S}^{yy}(\bm{k})$ and $\mathcal{S}^{zz}(\bm{k})$ at momentum $M_{2}$We observe the peak only at one of the nonequivalent $M$ wave vectors, since the $C_{3}$ rotational symmetry of the Hamiltonian is slightly broken by numerical rounding in the coupling constants obtained from the ab-initio calculations..
-As noticed previously [71], the symmetries of Eq.
-1 allow for two distinct stripe orders, differing by how they transform under the combined lattice-spin rotation around the triangular lattice axis directions.
-The first, we here call stripe-$\parallel$ (stripe-parallel), aligns the spins along the direction of the stripe.
-It is invariant under the $C_{2}$ axis rotations as the spins align with the axis of rotation.
-The stripe-$\perp$ (stripe-perpendicular) state aligns the spins perpendicular to the stripe direction, and thus transforms non-trivially under the $C_{2}$ axis rotation.
+The low energy spectra reveals an approximate GS degeneracy between the $\Gamma$.A, and $M.\text{A}$ irreps, consistent with the stripy AFM state [72].
+Accordingly, we observe a pronounced peak in $\mathcal{S}^{yy}(\bm{k})$ and $\mathcal{S}^{zz}(\bm{k})$ at momentum $M_{2}$The peak appears at only one $M$ point because slight numerical rounding in the ab-initio couplings weakly breaks the Hamiltonian’s $C_{3}$ symmetry..
+As noted [50], the symmetries of Eq.
+(1) allow for two distinct stripe orders, differing by how they transform under the combined lattice-spin rotation around the triangular lattice axis directions.
+The first aligns the spins along the direction of the stripe, we call it stripe-$\parallel$ (stripe-parallel) (this is also named stripe-$x$).
+It is invariant under the $C_{2}$ axis rotations, as the spins align with it.
+In the second, the spins align perpendicular to the stripe direction, and thus transforms non-trivially under the $C_{2}$ axis rotation.
+We name it stripe-$\perp$ (stripe-perpendicular) (also known as stripe-$yz$).
+Thus, the (in)variance under this symmetry constitutes the distinguishing feature between the two states .
 A representative configuration of each stripe type is shown in Fig.
-3d).
-In general, this state is not invariant under the $C_{2}$ axis rotation.
-Thus, the (in)variance under this symmetry constitutes the distinguishing feature between the two states.
-We prefer to refer to these states as stripe-$\parallel$ and stripe-$\perp$ as we think this more precisely captures the essence of these phases than their previously used names stripe-$x$ and stripe-$yz$, respectively.
-As the ground state stripe stabilized is oriented along the $x$-direction of the lattice, the peak in the $\mathcal{S}^{yy}(\bm{k})$ and $\mathcal{S}^{zz}(\bm{k})$ at $M_{2}$ indicates the realization of the stripe-$\perp$ state, see the SM for further details.
-Remarkably, this is in full agreement with experiments on KCeS2 [24].
+3 $d)$.
+The stabilized stripe is oriented along the $x$-direction of the lattice, so, the peak in the $\mathcal{S}^{yy}(\bm{k})$ at $M_{2}$ indicates the realization of the stripe-$\perp$ state, see the SM for further details.
+Remarkably, this is in complete agreement with experiments on KCeS2 [24].
 […]
-We obtain an approximate phase diagram, shown in panel $a)$ and $d)$ of Fig.
-3, based on the quantum numbers of the spectrum, which we show in the supplementary material.
-We begin by fixing $J^{\prime}=0.02J$ (a value numerically compatible with the largest matrix elements of the NNN coupling $\hat{J^{\prime}}$ in YZGO; see Table 3) and study the phase diagram as a function of the diagonal XY anisotropy $J_{\pm\pm}$ and the off-diagonal XY coupling $J_{z\pm}$.
+We perform extensive ED calculations on the $32$-site cluster and use the $12$-site cluster to corroborate the phase diagram obtained, whose results and discussion are left to the SM [63].
+We obtain an approximate phase diagram (Fig.
+3 $a)$ and $d)$) based on the quantum numbers of the spectrum (check the SM).
+We begin by fixing $J^{\prime}=0.02J$ (a value compatible with the largest matrix elements of the NNN coupling $\hat{J^{\prime}}$ in YZGO; see Table 2) and study the phase diagram as a function $J_{\pm\pm}$ and $J_{z\pm}$.
 This corresponds to the panels ($a,b,c$) in the upper row of Fig.
 3.
-In this case, the phase diagram consists of three distinct magnetically ordered states, namely, a $120^{\circ}$ AFM region, a stripe-$\parallel$ phase, and a stripe-$\perp$ phase, which correspond to the dark red, dark blue, and pale purple regions shown in panel $a)$, respectively.
-As on the 32-site cluster, the $K$ momentum is not resolved, the structure factor peaks at nearby wave vectors $Z_{0}$ and $Z_{1}$ in the $120^{\circ}$ AFM phase, as shown in panel (b) of Fig.
-3.
+In this case, the phase diagram consists of three distinct magnetically ordered states, a $120^{\circ}$ AFM region, a stripe-$\parallel$ phase, and a stripe-$\perp$ phase, which correspond to the dark red, dark blue, and pale purple regions shown in panel $a)$, respectively.
+As in the 32-site cluster, the $K$ momentum is not resolved, $\mathcal{S}(\bm{k})$ peaks at nearby wave vectors $Z_{0}$ and $Z_{1}$ in the $120^{\circ}$ AFM phase, as shown in Fig.
+3 $b)$.
 This reflects the frustrated realization of $120^{\circ}$ AFM order in this cluster.
-In contrast, the 12-site cluster resolves the $K$ points, showing a structure factor peak there consistent with the $120^{\circ}$ AFM order (see supplementary material).
-We observe in panel b) that the couplings determined for the YZGO system are within the $120^{\circ}$ AFM as expected.
-Upon tuning both $J_{\pm\pm}$ and $J_{z\pm}$, $\mathcal{S}(\bm{k})$ develops a peak at the $M$ points, signaling the onset of the stripe phase, as shown in panels $c)$ and $d)$.
-From the phase diagram in panel (a), we find that along the $J_{z\pm}=0$ line, negative values of $J_{\pm\pm}/J$ favor a stripe-$\parallel$ GS, whereas positive values of $J_{\pm\pm}/J$ stabilize a stripe-$\perp$ state.
+Upon tuning both $J_{\pm\pm}$ and $J_{z\pm}$, $\mathcal{S}(\bm{k})$ develops a peak at the $M$ points, signaling the onset of the stripe phase (see panel $c)$) We find that along the $J_{z\pm}=0$ line, negative values of $J_{\pm\pm}/J$ favor a stripe-$\parallel$, whereas positive values stabilize a stripe-$\perp$ state.
 The coupling $J_{z\pm}$ further extends the stripe-$\perp$ region into negative $J_{\pm\pm}/J$ values, which agrees with previous studies.
-In certain parameter regimes, the peak in $\mathcal{S}(\bm{k})$ appears in both the $M_{0}$ and $M_{1}$ wave vectors, while in others it shifts to the $M_{2}$ wave vector.
-This shift does not signify a transition between the two stripe phases, but instead arises from the geometry of the $32$-site cluster (being absent in the $12$-site cluster).
-As the $32$-site cluster does not have the $C_{3}$ rotation symmetry, the three $M$ points are inequivalent.
-As a result, the degeneracy among the six stripe configurations is lifted by the lattice geometry.
-To account for this, in panel (c) of Fig.
-3, we display
-$\tilde{\mathcal{S}}(M)=[\mathcal{S}(M_{0})+\mathcal{S}(M_{1})+\mathcal{S}(M_{2})]/3$ (4)
-averaged over the three $M$ points. Distinguishing between the two stripe types requires examining additional spin-spin correlators, which is left to the SM.
-We now set $J_{z\pm}=0$ and study the phase diagram as a function of the NNN coupling $J^{\prime}$ and $J_{\pm\pm}$.
-Based on the analysis of the quantum numbers of the lowest excited states, we map out the phase diagram identifying four distinct regions, as shown in panel $d)$ of Fig.
-3.
-For $J^{\prime}/J\sim[0.0,0.05]J$, the GS is in an $120^{\circ}$ AFM state, which is further corroborated by the peak in $\mathcal{S}(\bm{k})$ at the $Z$ points (shown in panel $e)$).
-The $120^{\circ}$ AFM is stabilized up to values of $J_{\pm\pm}\sim 0.1J$.
-Increasing $J_{\pm\pm}$ promotes the development stripe order, whose precise nature depends on the sign of $J_{\pm\pm}/J$, as shown in panel $d)$.
-Negative values favor a stripe-$\parallel$ GS, whereas positive values stabilize a stripe-$\perp$.
-The stripe order is further corroborated by the strong intensity of $\mathcal{S}(\bm{k})$ at the $M$ points, as shown in the panel $f)$.
-The black and white lines indicating the phase boundaries in Fig.
-3 are obtained from the position of level crossings in the excitation spectrum, see SM [56] for details.
-In the intermediate regime, $J^{\prime}\sim[0.05,0.16]J$ and $J_{\pm\pm}\sim\pm 0.1J$, the first excited state outside the $\Gamma.\text{A}$ sector transforms according to the $X_{1}.\text{A}$ irrep, as shown in the supplementary material.
-This parameter range corresponds to the orange region in the phase diagram (panel $d)$).
-In the context of triangular-lattice antiferromagnets, the appearance of an excitation at momentum $X$ is particularly noteworthy, as it has been interpreted as a signature of a DSL [19, 20].
-Specifically, Ref.
-[19] predicted that the singlet-monopole excitation of the DSL carries momentum $X$ and appears in the low-energy spectrum of the $J_{1}$$J_{2}$ triangular-lattice antiferromagnet in the spin-liquid regime [11].
+In certain parameter regimes, the peak in $\mathcal{S}(\bm{k})$ appears at both $M_{0}$ and $M_{1}$, while in others it shifts to $M_{2}$.
+This shift does not signal a transition between stripe phases, but originates from the absence of $C_{3}$ symmetry in the $32$-site cluster, which renders the three $M$ points inequivalent (unlike in the $12$-site cluster).
+Consequently, the lattice geometry lifts the degeneracy among the six stripe configurations.
+To account for this, in Fig.
+3 $c)$ we plot $\mathcal{S}(\bm{k})$ averaged over the three $M$ points,
+$\tilde{\mathcal{S}}(M)=\frac{\mathcal{S}(M_{0})+\mathcal{S}(M_{1})+\mathcal{S}(M_{2})}{3}.$ (3)
+Distinguishing the two stripe types requires examining additional spin-spin correlators, which we do in the SM.
+We observe in panel $b)$ that the couplings determined for the YZGO system place it within the $120^{\circ}$ AFM region as expected.
+We set $J_{z\pm}=0$ and study the phase diagram as a function of $J^{\prime}$ and $J_{\pm\pm}$, identifying four distinct regions, as shown in Fig.
+3 $d)$.
+For $J^{\prime}\in[0,0.05]J$, the GS is in a $120^{\circ}$ AFM state, confirmed by the peak of $\mathcal{S}(\bm{k})$ at the $Z$ points (panel $e)$).
+This state persists up to $J_{\pm\pm}\sim 0.1J$.
+Increasing $J_{\pm\pm}$ induces stripe order, whose orientation depends on the sign of $J_{\pm\pm}$: negative values yield stripe-$\parallel$, while positive values stripe-$\perp$ GS.
+This is corroborated by the peak of $\mathcal{S}(\bm{k})$ at the $M$ points as shown in panel $f)$.
+For $J^{\prime}\sim[0.05,0.16]J$ and $J_{\pm\pm}\sim[-0.1,0.1]J$, the first non-zero momentum excitation belongs to the $X_{1}.\text{A}$ irrep (see SM), corresponding to the orange region in panel $d)$.
+As discussed, the excitation at momentum $X$ is particularly noteworthy, as it is interpreted as a signature of the singlet-monopole of a DSL [19, 20], predicted to appear in the low-energy spectrum of the $J_{1}$$J_{2}$ triangular-lattice antiferromagnet in the spin-liquid regime [11].
 An alternative explanation for such a low-lying excitation is a 12-site valence-bond solid, for which consistent correlations have also been observed in the spin-liquid regime [11].
-In this parameter window, the static spin structure factor exhibits nearly the same intensity at the $Z$ and $M$ points.
-We further find that this region is stabilized only for small values of $J_{\pm\pm}\sim\pm 0.1J$, while larger values drive the system towards the stripe phase.
-Conclusions – We have derived from first principles the crystal-field splitting, Kramer’s GS and intersite exchange interactions within the GS manifold for a set of rare-earth triangular materials, including two systems (KCeO2 and KCeS2) that order magnetically, as well as, two systems (RbCeO2 and YZGO) that have been reported as putative QSLs.
-The calculated crystal-field splitting and gyromagnetic $g$-factors are in good agreement with experiment.
-The mean-field Curie-Weiss temperature evaluated from the resulting low-energy Hamiltonian including both the calculated intersite exchange and classical dipole-dipole magnetic interaction agrees well with experiment for all three oxide systems but is severely underestimated for KCeS2.
-Solving the resulting low-energy Hamiltonian with ED, we find evidence for the $120^{\circ}$ AFM order in all systems apart from KCeS2, where a stripe-$\perp$ AFM state is instead stabilized, consistent with experimental observations.
+We find that this region is stabilized only for small values of $J_{\pm\pm}\sim\pm 0.1J$, while larger values drive the system towards the stripe phase.
+This analysis indicates that the previously reported quantum paramagnetic phases in earlier DMRG studies [54] (identified even in the absence of NNN coupling) might not correspond to a DSL phase.
+In the SM, we provide additional results of the energy spectra, which reveal that the $X.\text{A}$ states persist at higher energies, whereas the low-lying excitations belong to the tower of states characteristic of the $120^{\circ}$ AFM order.
+While these results do not definitively exclude other paramagnetic states, the sharp $K$-point peak in the spin structure factor (computed also for an $N=36$ cluster at $J^{\prime}=0$, $J_{z\pm}=0.3J$, and $J_{\pm\pm}=-0.08J$) strongly supports $120^{\circ}$ AFM order.
+Although a DSL may exhibit similar quasi–long-range signatures, which finite-size ED cannot unambiguously distinguish from true long-range order, our spectral analysis offers little evidence for such a phase at this parameter point.
+Conclusions – We have derived from first principles the crystal-field splitting and and intersite exchange interactions within the Kramers GS manifold for a set of rare-earth triangular materials, including two systems (KCeO2 and KCeS2) that order magnetically, as well as two putative QSLs (RbCeO2 and YZGO).
+The calculated CF splitting and CF GS $g$-factors are in good agreement with experiment.
+The mean-field Curie-Weiss temperature evaluated from the resulting low-energy Hamiltonian agrees well with experiment for all three oxide systems but is severely underestimated for KCeS2.
+Solving the resulting low-energy Hamiltonian with ED, we find evidence for the $120^{\circ}$ order in all systems apart from KCeS2, where a stripe-$\perp$ state is instead stabilized, consistent with experimental observations.
 We find strong evidence for a magnetic $120^{\circ}$ order in the two DSL candidates.
@@
 However, in rare-earth semiconductors there can be another type of $f$-$f$ intersite exchange generated through 4ff-5dd hopping and the inter-shell Hund’s rule interaction [76].
-This could become important in the case of very weak superexchange like in KCeS2, where our calculations apparently reproduce the structure of intersite exchange but not its magnitude.
+This mechanism could become important in the case of very weak superexchange like in KCeS2, where our calculations apparently reproduce the structure of intersite exchange but not its magnitude.
+Its impact on the NNN exchange $J^{\prime}$ also needs to be evaluated, especially in the light of our present predictions for $J^{\prime}$ to be too small to stablize a DSL phase in YZGO.
 Further studies to evaluate the importance of such exchange mechanisms in triangular rare-earth compounds are worthwhile.
@@
 L.V.P.
-is grateful to Sidhartha Dash, Dario Fiore Mosca, and Antoine Georges for useful discussions and to the CPHT computer team for support.
+is grateful to Sidhartha Dash, Dario Fiore Mosca, and Antoine Georges for useful discussions.
+L.V.P acknowledges the support by CNRS through the Tremplin@Physique 2025 program and by the CPHT computer team.
 A.W.
@@
 509755282).
-Supplemental Material to ‘Ab initio spin Hamiltonians and magnetism of Ce and Yb triangular-lattice compounds’
-Appendix A Ab initio methodology and calculational details
-[…]
-[…]
-In the case of Ce delafossites, the Ce 4ff KS band manifold is separated from the ligand $p$-band by a wide gap and it only slightly overlaps with the Ce 5dd band.
-The lower boundary of the energy window is thus chosen to be within the $pf$ gap; the upper boundary is placed at 1.5 eV above the KS Fermi level $E_{F}^{KS}$, right above the top of the 4ff band.
-[…]
-In order to correct for this problem in YZGO, we thus apply, within Wien2k, a uniform shift to the 4ff band position to keep its centerweight position aligned to that in pure DFT (0.9 eV below $E_{F}^{KS}$) in the course of DFT+HI iterations.
-To disentangle 4ff from other contributions to hybridized 2pp-4ff bands, we extend the Wannier basis to include ligand O 2pp and itinerant Zn 3dd states.
-It has been shown [79] that employing a narrow energy window to construct 4ff Wannier orbitals is crucial to capture the hybridization contribution to CF.
-Since this narrow window excludes most of O 2pp and Zn 3dd states, we employ two energy windows.
-The 4ff Yb orbitals are formed from applying the initial projection to the bands in the range [-1.7:1.7] eV around the centerweight of KS 4ff band, which fully encloses this band.
-The O 2pp and Zn 3dd Wannier orbitals are formed from projection in the large window of [-9.5:1.5] eV around $E_{F}^{KS}$ that fully includes both corresponding band manifolds.
-The resulting full set of Wannier orbitals including Yb 4ff, O 2pp and Zn 3dd is then orthonormalized using the standard prescription of the projective construction approach [49] to obtain a proper Wannier basis.
-We parametrize the 4ff-shell rotationally-invariant on-site Coulomb repulsion by $U=F^{0}$ and the Hund’s rule coupling $J_{H}$ using the values U=U=6(8) eV and $J_{H}=$0.7(1.1) eV for the Ce(Yb) 4ff shells, respectively.
-The values of $U$ and $J_{H}$ together with the standard assumptions on the ratios of the Slater parameters ($F^{4}/F^{2}$=0.668, $F^{6}/F^{2}$=0.45 [80]) fully determine the vertex.
-The chosen $J_{H}$ values are consistent with experimental data on RE 3+ ions in insulators [81], the values of $U$ are within the generally accepted range and increasing from Ce to Yb in agreement with theoretical estimates for the RE series [82, 83].
-We employed the fully localized limit for the DMFT double-counting correction, which was calculated using the nominal $f^{1}$ ($f^{13}$) occupancy in the Ce (Yb) cases as shown to be appropriate for the DFT+HI method [43].
-To avoid the DFT self-interaction error impacting the CF parameters we employ averaging over the ground-state multiplet in our self-consistent calculations in accordance with Ref.
-[45].
-Namely, the Boltzmann weights for all levels within the ground-state multiplet $^{2}F_{5/2}$ ($^{2}F_{7/2}$) of the Ce (Yb) 4ff shell are averaged in the course of self-consistent DFT+HI calculations.
-Having converged DFT+HI calculations, we compute the intersite exchange interactions between the CF ground-state Kramer’s doublets using the FT-HI method.
-This method extract the low-energy exchange Hamiltonian from the paramagnetic electronic structure by evaluating the response of DFT+DMFT grand potential to small fluctuations around the paramagnetic state on two neighboring sites, see Ref.
-[42] for details.
-To form a proper (pseudo-)spin-1/2, the phases of the Kramer’s doublet states are chosen to have the proper symmetry under the time reversal $T$, $T|+1/2\rangle=|-1/2\rangle$.
-Appendix B DFT+Hubbard-I paramagnetic electronic structure
-In Fig.
-S1 we display the electronic structure of the paramagnetic phase in Ce delafossites (exemplified by KCeO2) and YZGO calculated by the DFT+Hubbard-I approach.
-In the case of KCeO2, the Ce 4ff lower Hubbard band is located in the wide semiconducting gap between O 2pp and Ce 5dd.
-In the case of the Yb compounds, the 4ff lower Hubbard band is within the O 2pp band and also overalaps with the filled Zn 3dd states, while the 4ff upper Hubbard band is located about 3.5 eV above the bottom of the conduction band (the latter is mainly formed by Yb 5dd states hybridized with O 2pp).
-Figure S1: DFT+Hubbard-I spectral functions of KCeO2 (top panel) and YbZn2GaO5 (bottom panel).
-Appendix C Crystal field levels and wavefunctions
-[…]
-Energy (meV) Wavefunction
-KCeO2
-0 $0.853|5/2;\mp 1/2\rangle\pm 0.470|5/2;\pm 5/2\rangle\pm 0.177|7/2;\mp 1/2\rangle-0.134|7/2;\mp 7/2\rangle+0.041|7/2;\pm 5/2\rangle$
-117 $0.993|5/2;\pm 3/2\rangle\mp 0.116|7/2;\pm 3/2\rangle+0.039|7/2;\mp 3/2\rangle$
-167 $0.828|5/2;\pm 5/2\rangle\mp 0.484|5/2;\mp 1/2\rangle+0.243|7/2;\mp 1/2\rangle\pm 0.145|7/2;\mp 7/2\rangle$
-KCeS2
-0 $0.895|5/2;\mp 1/2\rangle\pm 0.444|5/2;\pm 5/2\rangle-0.045|7/2;\mp 7/2\rangle$
-38 $0.998|5/2;\pm 3/2\rangle\mp 0.055|7/2;\pm 3/2\rangle+0.032|7/2;\mp 3/2\rangle$
-68 $0.890|5/2;\pm 5/2\rangle\mp 0.441|5/2;\mp 1/2\rangle+0.087|7/2;\mp 1/2\rangle\pm 0.074|7/2;\mp 7/2\rangle$
-RbCeO2
-0 $0.872|5/2;\mp 1/2\rangle\pm 0.439|5/2;\pm 5/2\rangle\pm 0.169|7/2;\mp 1/2\rangle-0.132|7/2;\mp 7/2\rangle+0.039|7/2;\pm 5/2\rangle$
-109 $0.993|5/2;\pm 3/2\rangle\mp 0.111|7/2;\pm 3/2\rangle+0.040|7/2;\mp 3/2\rangle$
-164 $0.851|5/2;\pm 5/2\rangle\mp 0.453|5/2;\mp 1/2\rangle+0.229|7/2;\mp 1/2\rangle\pm 0.131|7/2;\mp 7/2\rangle$
-YbZn2GaO5
-0 $0.725|7/2;\mp 7/2\rangle-0.501|7/2;\pm 5/2\rangle\mp 0.471|7/2;\mp 1/2\rangle$
-21 $0.831|7/2;\mp 5/2\rangle+0.555|7/2;\pm 7/2\rangle\pm 0.033|7/2;\pm 1/2\rangle$
-54 $|7/2;\pm 3/2\rangle$
-84 $0.881|7/2;\mp 1/2\rangle\pm 0.408|7/2;\mp 7/2\rangle\mp 0.238|7/2;\pm 5/2\rangle$
-Table 1: Calculated crystal-field energies and wavefunctions in all four compounds.
-Figure S2: Geometry of the simulation clusters with (a) $N=12$, (b) $N=32$, and (c) $N=36$ spins. The shaded area indicates the simulation box, with periodic boundary directions taken along the edges.
-Appendix D Simulation clusters used in the ED calculations
-The simulation clusters that we use in our calculations are the $N=12,32$ and $36$ sites shown in Fig.
-S2.
-The shaded region corresponds to the simulation box with periodic boundary conditions taken along each of the edges.
-The $N=12$ and $36$ site clusters are compatible with the full point-group symmetries of the infinite triangular lattice, the full dihedral group of order 12 ($D_{6}$), whereas the $N=32$ site cluster is not due to its specific geometry and its symmetries correspond to the full dihedral group of order 4 ($D_{2}$).
-As discussed in the main text, the $12$ and $36$ site clusters are compatible with all the symmetries of the anisotropic Hamiltonian, while the $32$ site is not.
-Appendix E Quantum Numbers of the symmetry broken states
-In the thermodynamic limit, the ground state of magnetically ordered systems spontaneously breaks the symmetries of the Hamiltonian.
-This breaking is reflected in the organization of quantum numbers in the low-energy spectrum, forming the so-called Anderson Tower of States.
-Here we predict their quantum numbers using group representation theory, following Ref.
-[67].
-We perform this analysis for two types of stripe-ordered states, stripe-$\parallel$ and stripe-$\perp$.
-The corresponding classical product states are shown in panels $a)$ and $b)$ of Fig.
-S3, respectively.
-The two stripe orders differ in how they transform under combined lattice-spin $C_{2}$ rotation around the lattice axis.
-As a result, the quantum numbers in panels $a)$ and $b)$ differ in their point-group irrep: the stripe-$\parallel$ states transform under the trivial irrep, while the stripe-$\perp$ states transform under the odd irrep.
-The multiplicities of the irreps in the degenerate ground-state manifold are calculated considering the point-group $D_{3d}$.
-However, in our ED calculations we only use translational and $C_{2}$ lattice rotation symmetries (ignoring spin rotations).
-Consequently, we do not resolve the quantum numbers of the entire symmetry group shown in panel $a)$ and $b)$.
-Under the used symmetries in the ED calculations, the two types of stripe order have identical signatures in the spectra, as shown in panel $c)$.
-Nevertheless, we can identify the transition between the two stripe types through a level crossing in the spectrum, specifically of the states that transform under the irreps with momentum $M$.
-We have also carried out this analysis for the $120^{\circ}$ AFM state, again using only translational and $C_{2}$ lattice rotation symmetries.
-The corresponding quantum numbers are shown in panel $d)$.
-Figure S3: Scheme of the six degenerate classical ground states for $a)$ stripes-$\parallel$ and $b)$ stripes-$\perp$, together with the multiplicities ($n$) of the corresponding irreducible representations under the full $D_{3d}$ point-group symmetry of the Hamiltonian.
-c)d)c)-d) Multiplicities of the irreducible representations for both stripe-$\parallel$ and stripe-$\perp$ (table $c)$) and for the $120^{\circ}$ AFM phase (table $d)$) when the Hamiltonian possesses only the $C_{2}$ lattice symmetry.
-Appendix F ED results effective Hamiltonians of KCeO2 and RbCeO2
-Figure S4: $a)$ Low-energy spectrum, $b)$ $\mathcal{S}(\bm{k})$ and $c)$ spin correlations $\langle\bm{S}_{\bm{r}_{j}}\cdot\bm{S}_{\bm{r}_{0}}\rangle$ in the GS of the $36$-site cluster effective magnetic Hamiltonians for the different compounds considered in this work (each row corresponds to a different compound).
-The spin correlations are computed with respect to the reference site $r_{0}$ marked by the green cross.
-Here, we discuss the low-energy spectra and GS correlations on the effective Hamiltonians derived for the RbCeO2 and KCeO2 materials (consult Eq.
-(1) and Tab.
-3).
-These results are shown in the third and fourth row of Fig.
-S4.
-For comparison, we show in the same figure the results for the YZGO and KCeS2 compounds that were previously discussed in the main text.
-Similarly to the YZGO material, we observe an organization of the low-energy excitations and spin correlations pointing to the $120^{\circ}$ AFM order.
-First, in the low-energy excitations, we observe that the GS lies in the $\Gamma.A$ irrep, with the first excited state corresponding to the $\Gamma.\text{B}$ being almost degenerate with a pair of excitations at momentum $K$, as shown in panel $a)$ for both RbCeO2 and KCeO2 materials.
-These excitations in this symmetry sectors are compatible with those expected for the $120^{\circ}$ AFM phase (see Fig.
-S3 d)).
-Secondly, $\mathcal{S}(\bm{k})$ peaks at the $K$ wave-vectors (shown in panels $b)$) .
-The real-space spin correlations [panels $c$] exhibit the sign-changing pattern characteristic of the three-sublattice structure of the $120^{\circ}$ order, reflecting the alternating spin orientations imposed by the triangular lattice geometry and frustrated antiferromagnetic interactions.
-This pattern is found in all compounds except KCeS2, which instead shows GS correlations characteristic of stripe order.
-Along the $\hat{x}$ direction the spins align ferromagnetically, while antiferromagnetic correlations remain along $\hat{y}$, consistent with the strong peak of $\mathcal{S}(\bm{k})$ at momentum $M_{2}$.
-To determine whether the system realizes the stripe-$\parallel$ or stripe-$\perp$ phase, we analyze $\mathcal{S}^{xx}(\bm{k})$, $\mathcal{S}^{yy}(\bm{k})$, and $\mathcal{S}^{zz}(\bm{k})$, shown in panels a)a)$c)$ of Fig.
-S5.
-All three components peak at $M_{2}$, confirming stripe order oriented along $100]\parallel x$.
-However, the much stronger correlations in $\mathcal{S}^{yy}(M_{2})$ and $\mathcal{S}^{zz}(M_{2})$ indicate that the GS realizes the stripe-$\perp$ configuration.
-Figure S5: a) $\mathcal{S}^{xx}(\bm{k})$, b) $\mathcal{S}^{yy}(\bm{k})$ and c) $\mathcal{S}^{zz}(\bm{k})$ in the 36-site cluster for the effective Hamiltonian GS of the KCeS2 compound.
-Appendix G Low-energy spectrum for the $N=32$ cluster and Further Results on the Spin Structure Factor
-Here, we present a detailed analysis of the low-energy spectrum of the effective magnetic Hamiltonian on the $N=32$ site cluster used in the main text to construct the extended phase diagram.
-In addition, we discuss the static spin structure factor, which allow us to distinguish between the stripe-$\parallel$ and stripe-$\perp$ phases.
-We begin with the spectra.
-Fig.
-S6 shows the dependence of the low-energy levels on different symmetry sectors as a function of the model parameters.
-Panel (a) displays the dependence on $J_{\pm\pm}$ at fixed $J_{z\pm}=0.173J$, taken from Fig.
-3(a), while panel (b) shows the dependence on $J^{\prime}$ at fixed $J_{\pm\pm}=-0.0488J$, taken from Fig.
-3(d).
-The different magnetic phases can be identified from the organization of the quantum numbers.
-The $120^{\circ}$ AFM region is signaled by a first excited state in the $\Gamma.\text{A}$ irrep together with two degenerate excited states in the $Z_{0}.\text{A}$ and $Z_{1}.\text{A}$ irreps.
-The transition to either stripe phase is marked by the closure of the gap in the $\Gamma.\text{A}$ irrep, corresponding to the phase boundaries in the main text.
-This criterion agrees remarkably well with the magnetic order inferred from the spin structure factors, as shown in the main text.
-Deep in the stripe phase, one also expects three degenerate states in the $M_{0}.\text{A}$, $M_{1}.\text{A}$, and $M_{2}.\text{A}$ irreps, becoming quasi-degenerate with the ground state in the $\Gamma.\text{A}$ irrep, as clearly seen in Fig.
-S6(a).
-The putative DSL regime is indicated by a first excited state with non-zero momentum $X$.
-This regime is shown in Fig.
-S6(b), where the $X_{1}.\text{A}$ level becomes the lowest excitation with a finite momentum as $J^{\prime}/J$ increases.
-We emphasize that this criterion alone does not establish the existence of the DSL phase nor determine its precise boundaries.
-We represent the transition to and from this region by a dashed line to emphasize this point.
-As discussed above, the two types of stripe order cannot be distinguished solely from the quantum numbers of the spectra within the symmetries accessible to our ED calculations.
-However, they can be differentiated by analyzing the spin structure factor with spin components projected along the stripe direction and rotated by $90^{\circ}$ from it.
-Due to the symmetries of the $32$-site cluster, the six stripe configurations sketched in Fig.
-S3 are no longer degenerate.
-We find that in some regions of the phase diagram the ground state exhibits a characteristic wave vector at the $M_{2}$ point, while in others the peak appears at both $M_{0}$ and $M_{1}$ wave vectors.
-This distinction does not yet separate stripe-$\parallel$ from stripe-$\perp$; rather, it indicates which specific stripe arrangement (within each stripe type) is energetically preferred for this cluster geometry.
-So for the parameters where the order is characterized by the $M_{2}$ wave vector, we inspect $\mathcal{S}^{xx}(\text{M}_{2})$ and $\mathcal{S}^{yy}(\text{M}_{2})$, shown in Figs.
-S6(c) and (d).
-In the stripe-$\parallel$ phase, the spins preferably align along the stripe direction (the $x$-axis), leading to $\mathcal{S}^{xx}(M_{2})>\mathcal{S}^{yy}(M_{2})$.
-In contrast, in the stripe-$\perp$ phase the spins are rotated by $90^{\circ}$, resulting in $\mathcal{S}^{yy}(M_{2})>\mathcal{S}^{xx}(M_{2})$.
-From this criterion, we identify the stripe-$\parallel$ region for $J_{\pm\pm}>0$ and the stripe-$\perp$ region for $J_{\pm\pm}<0$.
-Nevertheless, there are regions within the stripe-ordered phases where neither $\mathcal{S}^{xx}(M_{2})$ nor $\mathcal{S}^{yy}(M_{2})$ show a peak.
-These correspond to regions where the GS develops stripe order with characteristic wave vectors at $M_{1}$ (or $M_{0}$).
-In this case, we analyze instead
-$\mathcal{S}^{\parallel\bm{a}_{2}}\left(\text{M}_{1}\right)=\dfrac{1}{N}\sum_{i,j}\left\langle\left(\bm{S}_{i}\cdot\hat{\bm{a}}_{2}\right)\left(\bm{S}_{j}\cdot\hat{\bm{a}}_{2}\right)\right\rangle,\quad\mathcal{S}^{\perp\bm{a}_{2}}\left(\text{M}_{1}\right)=\dfrac{1}{N}\sum_{i,j}\left\langle\left(\bm{S}_{i}\cdot[1-\hat{\bm{a}}_{2}]\right)\left(\bm{S}_{j}\cdot[1-\hat{\bm{a}}_{2}]\right)\right\rangle,$ (S1)
-where $\hat{\bm{a}}_{2}$ is the unit vector in the $\bm{a}_{2}$ lattice direction.
-These are shown in Figs.
-S6(e) and (f) as a function $J_{\pm\pm}/J$ and $J^{\prime}/J$.
-From these we obtain the same result: stripe-$\parallel$ is stabilized for $J_{\pm\pm}>0$, while stripe-$\perp$ is stabilized for $J_{\pm\pm}<0$.
-Moreover, the transition between them, occurring around $J_{\pm\pm}=0$, agrees remarkably well with the one identified from the crossing of the ground state and first excited state in the $M$-sector irreps.
-The behavior is identical when analyzing the dependence on $J_{\pm\pm}$ and $J_{z\pm}$ with $J^{\prime}$ fixed.
-Figure S6: Low-energy spectra and static spin structure factors of the effective magnetic Hamiltonian for the 32-site cluster.
-$a)$ - Spectrum as a function of $J_{\pm\pm}/J$ for $J^{\prime}=0.022J$ and $J_{z\pm}=0.173J$.
-$b)$- Spectrum as a function of $J^{\prime}/J$ for $J_{\pm\pm}=-0.0488J$ and $J_{z\pm}=0$.
-The colored regions indicate phases: blue – stripe-$\parallel$, dark blue – stripe-$\perp$, red – $120^{\circ}$ AFM, pale red – putative DSL.
-c)f)c)–f) - Static spin structure factors: (c) along $x$ and (d) along $y$ at momenta $M_{2}$; (e) parallel and (f) perpendicular to $\bm{a}_{2}$ at momenta $M_{1}$.
-Appendix H Low-energy spectrum and ground-state correlations for the $N=12$ cluster
-In this section, we discuss the phase diagram obtained for the 12-site cluster, whose results support those presented in the main text for the 32-site cluster.
-In contrast to the 32-site cluster, the 12-site cluster resolves the natural ordering vectors for both the stripe order and the $120^{\circ}$ AFM order, specifically the $M$ and $K$ wave-vectors, as depicted in Fig.
-S7 $d)$.
-Moreover, the $X$ points are also resolved, which is important for the DSL phase.
-Figure S7: a)-Approximate phase diagram as a function of $J_{\pm\pm}/J$ and $J_{z\pm}/J$ for different values of $J^{\prime}/J$ [i)–iii)].
-The blue cross marks the approximate coupling values of some of the compounds considered in this work (we are using the absolute value of $J_{z\pm}$ see the main text).
-$\mathcal{S}(\bm{k})$ is shown at the ordering vectors $K_{0}$ [iv–vi] and $M$ [vii–ix], as well as for the $x$ [x–xii] and $y$ [xiii–xv] spin components at the $M_{2}$ wave-vector.
-(b) Low-energy spectra resolved by symmetry sectors as a function of $J_{\pm\pm}/J$ for $J_{z\pm}=0.20J$ [i)] and $J_{z\pm}=0.69J$ [ii)], with $J^{\prime}=0.02J$.
-(c) Sketch of the FBZ for the 12-site cluster.
-Figure S8: $a)$ Approximate phase diagram based on the structure of the quantum numbers of the low-energy spectra as a function of $J_{\pm\pm}/J$ and $J^{\prime}/J$.
-The transition line between the two distinct stripe orders is obtained from the level crossing in the $M.\mathrm{A}$ irreps.
-$b)$ and $c)$ Low-energy spectra for $J_{\pm\pm}=0.0202J$ as a function of $J^{\prime}/J$ $b)$ and for $J^{\prime}=0.1054J$ as a function of $J_{\pm\pm}/J$ $c)$.
-The color scheme labeling the different irreps is the same as in Fig.
-S7.
-d)d)$f)$ Static spin structure factor for different as a function of $J_{\pm\pm}/J$ and $J^{\prime}/J$ evaluated at momentum $K$ (panel $d)$) and $M_{2}$ (panels $e)$ and $f)$).
-Other parameters: $N=12$ and $J_{z\pm}=0$.
-We begin by studying the phase diagram as a function of $J_{\pm\pm}/J$ and $J_{z\pm}/J$ for fixed values of $J^{\prime}$ consistent with those obtained for the compounds considered in this work (see Tab.
-3 in the main text).
-The phase diagram is shown in panel a), sub-panels i), ii), and iii), corresponding to $J^{\prime}=0.020J$, $J^{\prime}=-0.041J$, and $J^{\prime}=-0.062J$, respectively.
-The approximate phase boundaries are estimated from the quantum numbers of the low-energy spectrum.
-In the $120^{\circ}$ phase, the GS is unique transforming under the $\Gamma.\text{A}$ irrep, with low-energy excitations in the $\Gamma.\text{B}$, $K_{0}.\text{A}$ and $K_{1}.\text{A}$ irreps.
-We also observe an exactly degenerate first excited state also transforming under $\Gamma.\text{A}$ irrep.
-These exact degeneracy is due to the richer symmetry group than the one considered in our ED code.
-Upon tuning either $J_{\pm\pm}/J$ or $J_{z\pm}/J$, we observe a closure of the energy gap between the GS energy and the first excited state within the $\Gamma.{\rm A}$ irrep.
-The exact point where this gap vanishes is identified as the transition into the stripe phase.
-This transition is further confirmed by the decrease in intensity in $\mathcal{S}(K)$ after the gap closes, as shown in Fig.
-S7 a), panels iv)–vi).
-In the phase diagrams shown in panel a), we have located some of the compounds considered in this work.
-Their positions are determined by the calculated values of $^{\prime}/J, J±±/J$$_{\pm\pm}/J$, and the absolute value of $J_{z\pm}$ (magnetic phases with $J_{z\pm}<0$ can be mapped to those with $J_{z\pm}>0$ by applying a $\pi$-rotation to all spins in the system).
-We find their locations to be consistent with previous calculations, namely within the $120^{\circ}$ AFM state.
-In the stripe-ordered phase, the ground state becomes doubly degenerate in $\Gamma.{\rm A}$, while the first excited states are degenerate and transform under the $M_{0}.{\rm A}$, $M_{1}.{\rm A}$, and $M_{2}.{\rm A}$ irreps.
-This also corroborates the onset of stripy magnetic order, as these quantum numbers match those predicted from the representation group theory calculations (see panel c) in Fig.
-S3).
-This is again corroborated by the static spin structure factors evaluated at the $M$ point, as shown in Fig.S7 a), panels vii)–ix).
-To distinguish between the stripe-$\parallel$ and stripe-$\perp$,we examine $\mathcal{S}^{xx}(\bm{k})$ and $\mathcal{S}^{yy}(\bm{k})$.
-Both exhibit peaks at the wave vector $M_{2}$, but in different regions of the phase diagram: the $xx$ peak identifies the stripe-$\parallel$ phase, while the $yy$ peak signals the stripe-$\perp$ phase.
-The phase diagram obtained from them is consistent with the transition line inferred from the energy-level crossing at the $M$ points.
-This crossing reflects the transition between the two stripe orders, which carry distinct quantum numbers in the low-energy sector (although both remain located at momentum $M$).
-This crossing is clearly illustrated in Fig.
-S7c), panels i) and ii), where we show the energy spectra as functions of $J_{\pm\pm}/J$ for two representative values of $J_{z\pm}$: $J_{z\pm}\simeq 0.20J$ (i)) and $J_{z\pm}\simeq 0.69J$ (ii)).
-We find that the $120^{\circ}$ order is suppressed as either $J_{\pm\pm}$ or $J_{z\pm}$ is increased, with negative values of $J^{\prime}/J$ enlarging this region in the phase diagram.
-By tuning $J_{\pm\pm}$ and/or $J_{z\pm}$, the GS instead exhibits stripe order.
-Similar to the 32-site cluster, we observe that for $J_{z\pm}=0$, negative $J_{\pm\pm}$ values favors the stripe-$\parallel$ phase, while positive values stabilize the stripe-$\perp$ phase.
-The stripe-$\perp$ phase is also supported for negative values of $J_{\pm\pm}/J$, when $J_{z\pm}\neq 0$.
-For all these parameters, we do not find the first excited state (with a non-zero momentum) in the $X.\mathrm{A}$ irrep, which therefore does not indicate the presence of a DSL phase.
-In Fig.
-S8, we set $J_{z\pm}=0$ and study the phase diagram as a function of $J_{\pm\pm}/J$ and $J^{\prime}.J$.
-In panel (a), the phase boundaries are determined in the same way as for Fig.
-S7.
-For $J^{\prime}/J\lesssim 1\%$, the behavior matches that discussed previously: a $120^{\circ}$ AFM phase is stabilized for finite $J_{\pm\pm}/J$, which, upon increasing $|J_{\pm\pm}/J|$, gives way to a stripy phase.
-The nature of this phase depends on the sign of $J_{\pm\pm}/J$: for negative values it is stripe-$\parallel$, while for positive values it is stripe-$\perp$.
-This is corroborated by the static spin structure factors shown in panels d) to f): in the $120^{\circ}$ phase we find strong peaks at both $K$ points, whereas in the stripe-$\parallel$ phase the $S^{xx}$ component exhibits a pronounced signal along the $M_{2}$ vector, and in the stripe-$\perp$ phase the same wave-vector is dominant in the $S^{yy}$ component.
-The transition between the two stripe phases is again identified by locating the level crossing between the ground state and the first excited state within the $M.\text{A}$ irreps.
-Increasing the strength of NNN coupling, $J^{\prime}/J$, leads to the appearance of first excitated state with non-zero momentum at the the $X_{1}.A$ and $X_{2}.A$ irrep.
-This is clearly seen in panels (b) and (c), where we plot the low-energy spectrum as a function of $J^{\prime}$ for $J_{\pm\pm}=0.02J$ and as a function of $J_{\pm\pm}$ along the line $J^{\prime}=0.1054J$, respectively.
-As discussed in the main text, this behavior pins the possible DSL region.
-From panel (a), we observe that the range of $J_{\pm\pm}/J$ supporting this region shrinks as $J^{\prime}$ increases.
-For sufficiently large values of either coupling, the system transitions into a stripy ordered phase whose orientation again depends on the sign of $J_{\pm\pm}/J$ (negative for stripe-$\parallel$, positive for stripe-$\perp$).
-Exactly at $J_{\pm\pm}=0$, the model recovers the SU(2) and U(1) spin-symmetries.
-These findings are consistent with the results for the 32-site cluster presented in the main text: while a possible DSL regime is identified, all ab initio estimates of $J^{\prime}/J$ lie below the required range.